### v1

# Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation

Abstract
Model merging integrates the parameters of multiple models into a unified model, combining their diverse capabilities. Existing model merging methods are often constrained by fixed parameter merging ratios. In this study, we propose Mixup Model Merge (M3), an innovative approach inspired by the Mixup data augmentation technique. This method merges the parameters of two large language models (LLMs) by randomly generating linear interpolation ratios, allowing for a more flexible and comprehensive exploration of the parameter space. Extensive experiments demonstrate the superiority of our proposed M3 method in merging fine-tuned LLMs: (1) it significantly improves performance across multiple tasks, (2) it enhances LLMs’ out-of-distribution (OOD) robustness and adversarial robustness, (3) it achieves superior results when combined with sparsification techniques such as DARE, and (4) it offers a simple yet efficient solution that does not require additional computational resources. In conclusion, M3 is a simple yet effective model merging method that significantly enhances the performance of the merged model by randomly generating contribution ratios for two fine-tuned LLMs. The code is available at https://github.com/MLGroupJLU/MixupModelMerge.
Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation
Yue Zhou1  Yi Chang1,2,3  Yuan Wu1111Corresponding authors 1School of Artificial Intelligence, Jilin University 2Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 3International Center of Future Science, Jilin University yuezhou24@mails.jlu.edu.cn, yichang@jlu.edu.cn, yuanwu@jlu.edu.cn
1 Introduction
In the field of Natural Language Processing (NLP), the emergence of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; OpenAI, 2023; Chowdhery et al., 2023) represents a revolutionary breakthrough. With their remarkable capabilities, these models have demonstrated outstanding performance across various tasks (Jiao et al., 2023; Chang et al., 2024b; Nam et al., 2024; Xing, 2024; Guo et al., 2024), significantly advancing NLP technologies.
(a) Performance of the merged models obtained through Task Arithmetic and TIES-Merging with/without M3.
(b) (Left) Performance of merged models with/without M3 on OOD datasets. (Right) Adversarial robustness (PDR) of merged models with/without M3.
(c) The performance trend of the merged models obtained through the following methods: TIES-Merging without M3 and DARE, TIES-Merging with DARE, TIES-Merging with M3, and TIES-Merging with both M3 and DARE.
Figure 1: (a) M3 significantly boosts the model merging performance. (b) OOD robustness, and adversarial robustness of the merged models. (c) Combined with DARE, M3 delivers even better results. LM, MATH, and Code refer to the WizardLM13B, WizardMath-13B, and llama-2-13b-code-alpaca models. TA stands for Task Arithmetic and TIES stands for TIES-Merging.
Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al., 2021; Ding et al., 2023; Xia et al., 2024). However, SFT requires substantial computational resources and long training times (Brown et al., 2020; Chang et al., 2024a). To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al., 2024; Akiba et al., 2024). It effectively reduces the resource-intensive demands of SFT while preserving and even enhancing model performance.
A simple analogy for model merging is the Super Mario game, where the protagonist gains special abilities by absorbing power-up items. Similarly, merging model parameters integrates the strengths of different models, enabling more effective multi-task learning (Yu et al., 2024). However, existing model merging methods have some limitations, these methods heavily rely on predefined or heuristic parameter fusion strategies (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024) such that they fail to fully explore the parameter space, thereby restricting the potential of the merged model.
To address this issue, we draw inspiration from Mixup (Zhang, 2017) and propose a novel technique called Mixup Model Merge (M3). This method introduces randomness by dynamically adjusting the contribution ratios between models, making the model merging process more flexible and enabling a thorough exploration of the parameter space. M3 further unlocks the potential of model merging, significantly enhancing the generalization performance of the merged model while improving its out-of-distribution (OOD) and adversarial robustness (Wang et al., 2023; Zhu et al., 2023). As shown in Figure 2, the implementation of M3 is similar to the process of proportionally mixing magical potions in Harry Potter. Specifically, this method dynamically controls the parameter fusion ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$. By adjusting the parameter $\alpha$, we can precisely control the distribution of $\lambda_{m}$, allowing M3 to explore the parameter space of model merging more deeply and thus fully unleash the potential of the models, leading to improved overall performance.
We conducted extensive experiments with three fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively. Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang, 2017), we further performed comprehensive evaluations on LiveBench (White et al., 2024) and PromptBench (Zhu et al., 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness of merged models. The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)), and boost model merging when combined with sparsification techniques like DARE (Yu et al., 2024) (as shown in Figure 1(c)).
Figure 2: Implementation of m3: A process analogous to proportionally mixing magical potions in harry potter. The proposed method controls the contribution ratio between two fine-tuned llms by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$. The distribution of $\lambda_{m}$ is controlled by adjusting $\alpha$.
2 Related Works
Model Merging
Model merging is a technique that integrates the parameters of multiple models to create a unified model with enhanced or diverse capabilities (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024; Lin et al., 2024). Task arithmetic (Ilharco et al., 2022) leverages task vectors for model merging through arithmetic operations, incorporating a predefined scaling term to weight the contribution of different models. Fisher Merging (Matena and Raffel, 2022) performs parameter fusion by applying weights derived from the Fisher information matrix (Fisher, 1922), resulting in more precise parameter integration. TIES-Merging (Yadav et al., 2023) addresses task conflicts by removing low-magnitude parameters, resolving sign disagreements, and merging only the parameters that align with the final agreed-upon sign. In (Yu et al., 2024), it is found that LLMs can enhance their capabilities through model merging. Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al., 2022), significantly improving the performance of various model merging techniques.
Mixup
Mixup is proposed to enhance the generalization ability of deep learning models by surpassing traditional Empirical Risk Minimization (ERM) (Zhang, 2017). It is a simple, data-agnostic augmentation technique that trains models using virtual examples created by linearly interpolating pairs of random examples and their corresponding labels. Rooted in the Vicinal Risk Minimization (VRM) principle (Chapelle et al., 2000), this approach improves generalization across a variety of datasets (Russakovsky et al., 2015; Krizhevsky et al., 2009; Warden, 2017; Asuncion et al., 2007), and helps reduce overfitting, sensitivity to adversarial examples, and training instability, all with minimal computational cost. Given two samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$, Mixup generates a new sample using the following formulas:
[…]
$\tilde{x}$ and $\tilde{y}$ represent the generated synthetic sample and its label, respectively, with $\lambda$ determining their interpolation ratio, typically ranging from 0 to 1. Here, $\lambda$ is a hyperparameter sampled from a Beta distribution, i.e., $\lambda\sim\text{Beta}(\alpha,\alpha)$, where $\alpha$ controls the strength of the interpolation between feature-target pairs. Inspired by Mixup, we propose a novel model merging method, M3, which generates the parameters of the merged model by performing linear interpolation between the parameters of two fine-tuned LLMs, with the interpolation ratio being random.
3 Methodology
3.1 Model Merging Problem
Following Yu et al. (2024), we focus on merging fine-tuned LLMs that have been optimized from the same pre-trained backbone (Touvron et al., 2023). Specifically, we aim to fuse the parameters of these LLMs to create a unified model capable of handling multiple tasks. In this context, we restrict our attention to the merging of two models, as the mixup theory, which forms the basis of our approach, is generally applied to two entities.
Given two tasks $t_{1}$ and $t_{2}$ with the corresponding fine-tuned LLMs having parameters $\theta_{\text{SFT}}^{t_{1}}$ and $\theta_{\text{SFT}}^{t_{2}}$, model merging aims to combine the parameters of these two models into a single model with parameters $\theta_{M}$. The resulting model should be able to effectively perform both tasks simultaneously, leveraging the knowledge learned from each individual model.
3.2 Mixup Model Merge
Inspired by Mixup (Zhang, 2017), we propose a simple yet effective model merging method called Mixup Model Merge (M3). Unlike existing methods that use fixed merging ratios, M3 generates the model contribution ratio randomly, harnessing randomness to inject fresh vitality into the model merging process. Specifically, M3 further explores the parameter space of the merged model to unlock the potential of model merging.
To further elaborate, given two fine-tuned LLMs with parameters $\{\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}}\}$, we combine M3 with established model merging methods to fuse these parameters and obtain a single merged model with parameters $\theta_{M}$. As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022).
The official computation process for Average Merging is described as follows:
$\theta_{M}=\frac{1}{2}\left(\theta_{\text{SFT}}^{t_{1}}+\theta_{\text{SFT}}^{t_{2}}\right)$ (2)
The official computation process for Task Arithmetic is:
$\theta_{M}$ $=\theta_{\text{PRE}}+\lambda\cdot(\delta^{t_{1}}+\delta^{t_{2}})$
$=\theta_{\text{PRE}}+\lambda\cdot\sum_{i=1}^{2}(\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}}),$ (3)
where $\theta_{\text{PRE}}\in\mathbb{R}^{d}$ represents the parameters of the pre-trained language model (PLM), such as Llama 2 (Touvron et al., 2023). $\lambda$ is a scaling factor that weights the contribution of each model during the merging process. $\delta^{t_{i}}$ denotes the delta parameter (Ilharco et al., 2022), which is defined as the difference between the parameters of the language models (LMs) before and after SFT, i.e., $\delta^{t}=\theta_{\text{SFT}}^{t}-\theta_{\text{PRE}}\in\mathbb{R}^{d}$, where $t$ refers to task $t$.
When introducing M3, the process for Average Merging is reformulated as:
$\theta_{M}=\lambda_{m}\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\theta_{\text{SFT}}^{t_{2}},$ (4)
while the process for Task Arithmetic is reformulated as:
$\theta_{M}$ $=\theta_{\text{PRE}}+\lambda_{m}\delta^{t_{1}}+(1-\lambda_{m})\delta^{t_{2}},$ (5)
where $\lambda_{m}$ determines the linear interpolation ratio between the two fine-tuned LLMs, and is generally a value between 0 and 1. $\lambda_{m}$ is sampled from a Beta distribution, typically $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$, where $\alpha$ controls the shape of the Beta distribution.
The hyperparameter $\alpha$ for M3 is selected from the range $[0.2,0.4,0.5,1,2,3,5]$. As shown in Figure 3: (1) When $\alpha=1$, $\lambda$ follows a uniform distribution, meaning all values within the range $(0,1)$ are equally likely to be sampled. (2) When $\alpha<1$, the distribution of $\lambda$ exhibits a bimodal shape, with higher probabilities near the extremes (0 and 1). This indicates that the merged model is more likely to be dominated by one of the two models. (3) When $\alpha>1$, the distribution of $\lambda$ becomes concentrated around the middle (e.g., 0.5), resulting in more balanced contributions from both models.
Figure 3: The Beta distribution visualization for different $\alpha$ values.
3.3 Theoretical Analysis
Mixup performs linear interpolations between data samples and their labels in the data space (Zhang, 2017). Similarly, M3 can be viewed as applying random linear interpolation in the parameter space, where the interpolation occurs between the parameters of two fine-tuned models trained on different tasks. M3 represents a natural extension of the Vicinal Risk Minimization (VRM) principle (Chapelle et al., 2000) into the model parameter space. In data space, VRM introduces a vicinal distribution to simulate the true data distribution, thereby increasing the diversity of training data and enhancing the model’s generalization ability. Similarly, M3 extends this principle by constructing a virtual neighborhood in the model parameter space between two task-specific models, combining their knowledge in a way that is more natural and balanced.
In the context of model merging, interpolating between two sets of model parameters, $\theta_{\text{SFT}}^{t_{1}}$and $\theta_{\text{SFT}}^{t_{2}}$, defines a new neighborhood distribution in the parameter space, which can be expressed as:
[…]
where $P_{\nu}(\theta_{M})$ represents the probability distribution of the new model parameters $\theta_{M}$ generated through interpolation. The function $\nu(\theta_{M}\mid\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}})$ defines the range and behavior of $\theta_{M}$, depending on the interpolation strategy (such as linear interpolation). Additionally, $P(\theta_{\text{SFT}}^{t_{1}})$ and $P(\theta_{\text{SFT}}^{t_{2}})$ represent the prior distributions of the parameters of the two models, respectively.
Under the linear interpolation strategy, the function $\nu(\theta_{M}\mid\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}})$ is defined as:
$\theta_{M}=\lambda_{m}\cdot\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\cdot\theta_{\text{SFT}}^{t_{2}},$ (7)
where $\lambda_{m}\in(0,1)$ is the interpolation ratio. The corresponding distribution $P_{\nu}(\theta_{M})$ is expressed as:
where $\delta(\cdot)$ is the Dirac delta function, ensuring that $\theta_{M}$ satisfies the linear interpolation rule. By using this linear interpolation method, the parameters $\theta_{\text{SFT}}^{t_{1}}$ and $\theta_{\text{SFT}}^{t_{2}}$ are combined in the parameter space, forming a new neighborhood distribution $P_{\nu}(\theta_{M})$. This process effectively merges the knowledge of both models into a unified parameter space, thus achieving robust performance across different tasks while maintaining the original strengths of the models.
By interpolating the parameters, M3 encourages the model to learn smoother decision boundaries. In this context, smoother decision boundaries can be understood as the boundaries between different tasks, such as instruction following, mathematical reasoning, and code generation, where the model must understand and adapt to each task differently. In tasks $t_{1}$ and $t_{2}$, M3 creates a virtual neighborhood that seamlessly integrates knowledge from both tasks. This approach prevents the merged model from overfitting task-specific details, ensuring a balanced and effective performance across all tasks.
Secondly, M3 introduces a linear inductive bias in the parameter space, encouraging the merged model parameters to lie on a linear manifold between the two source models. This linear structure offers significant advantages in terms of simplicity and generalization. According to Occam’s Razor, simpler solutions tend to generalize better. By performing linear interpolation between two sets of parameters, M3 avoids unnecessary complexity in the model merging process, leading to a more straightforward and efficient solution.
Thirdly, M3 can improve the performance of the merged model across multiple tasks by mitigating task conflicts. Different tasks may require conflicting parameter values, which can lead to performance degradation on one task while optimizing for another. Linear interpolation helps balance these conflicts, resulting in a model that performs well among all tasks.
4 Experiments
4.1 Experimental Setup
Task-Specific Fine-Tuned LLMs and Datasets
Following the experimental setup given in Yu et al. (2024), we select three task-specific fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), all of which use Llama-2-13b (Touvron et al., 2023) as the pre-trained backbone. These models are respectively designed for instruction-following, mathematical reasoning, and code generation tasks. To evaluate the instruction-following task we use AlpacaEval (Li et al., 2023). For testing mathematical reasoning task, we employ GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For estimating the performance of code-generating task, we use HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). More details of these LLMs and datasets can be found in Appendix A.1.
The Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness
To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al., 2024), respectively. More details on OOD benchmarks are given in Appendix A.3.
We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al., 2024) to assess the robustness of LLMs against adversarial prompts. Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al., 2018), BERTAttack (word-level) (Li et al., 2020), and StressTest (sentence-level) (Naik et al., 2018). The evaluation is conducted on two datasets supported by PromptBench: SST2 (sentiment analysis) (Socher et al., 2013) and CoLA (grammatical correctness) (Warstadt, 2019). For more details on PromptBench and attack methods, please refer to Appendix A.4.
Evaluation Metrics
We calculate win rate for AlpacaEval and LiveBench-Instruction, zero-shot accuracy for GSM8K and MATH, pass@1 for HumanEval, MBPP and LiveBench-Coding, Matthews correlation coefficient (MCC) for CoLA, accuracy for SST2, and zero-shot accuracy for LiveBench-TypoFixing.
Implementation Details
Unless otherwise specified, the details of the model merging experiments are consistent with Yu et al. (2024). The hyperparameter $\alpha$ for M3 is chosen from the range $[0.2,0.4,0.5,1,2,3,5]$. For a detailed description of the hyperparameter settings in model merging methods, please refer to Appendix A.2. Additionally, all experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
4.2 Merging Task-Specific Fine-Tuned LLMs
Merging Methods Models Use Model Mixup Use DARE Instruction Following Mathematical Reasoning Code Generating
AlpacaEval GSM8K MATH HumanEval MBPP
/ LM No No 45.14 2.20 0.04 36.59 34.00
Math No No / 64.22 14.02 / /
Code No No / / / 23.78 27.60
Average Merging LM & Math No No 45.28 66.34 13.40 / /
Yes No 44.40 66.26 13.80 / /
No Yes 44.22 66.57 12.96 / /
Yes Yes 43.53 66.57 14.12 / /
LM & Code No No 36.60 / / 29.88 32.00
Yes No 43.91 / / 37.20 34.40
No Yes 38.81 / / 31.71 32.40
Yes Yes 40.31 / / 36.59 37.00
Math & Code No No / 56.17 10.28 8.53 8.20
Yes No / 63.61 14.02 8.54 19.20
No Yes / 56.18 10.28 6.10 7.80
Yes Yes / 64.97 13.54 9.76 21.20
Task Arithmetic LM & Math No No 45.78 66.34 13.40 / /
Yes No 41.65 66.34 13.74 / /
No Yes 49.00 66.64 13.02 / /
Yes Yes 44.90 67.32 13.74 / /
LM & Code No No 44.64 / / 32.93 33.60
Yes No 46.64 / / 35.37 33.80
No Yes 41.47 / / 35.98 33.00
Yes Yes 45.20 / / 35.98 35.20
Math & Code No No / 64.67 13.98 8.54 8.60
Yes No / 63.53 13.94 7.93 19.00
No Yes / 65.05 13.96 10.37 9.80
Yes Yes / 65.13 14.32 8.54 18.00
TIES- Merging LM & Math No No 38.63 14.56 2.12 / /
Yes No 38.73 18.57 2.48 / /
No Yes 37.92 18.04 2.34 / /
Yes Yes 39.93 19.26 2.82 / /
LM & Code No No 41.85 / / 0.0 0.0
Yes No 44.96 / / 25.61 30.80
No Yes 43.13 / / 0.0 0.0
Yes Yes 45.65 / / 26.83 33.20
Math & Code No No / 64.67 13.68 9.15 22.60
Yes No / 64.75 14.16 9.76 21.4
No Yes / 64.82 13.88 10.37 23.60
Yes Yes / 64.75 14.78 9.15 19.60
Table 1: Performance of merging task-specific LLMs WizardLM-13B (LM), WizardMath-13B (Math), and llama-2-13b-codealpaca (Code) on all the datasets. The best and second-best results are marked in bold and underlined fonts.
We integrate M3 into three prominent model merging techniques: Average Merging, Task Arithmetic, and TIES-Merging. The performance of merging task-specific fine-tuned LLMs is presented in Table 1.
From Table 1, we obtain the following observations: 1) M3 generally enhances Average Merging, Task Arithmetic, and TIES-Merging when merging fine-tuned LLMs. For example, the improvements achieved by Average Merging with M3 for Math & Code are 7.43% on GSM8K, 3.74% on Math, and 11.0% on MBPP. For LM & Code, Average Merging with M3 shows improvements of 7.31% on AlpacaEval, 7.32% on HumanEval, and 2.4% on MBPP. Task Arithmetic with M3 results in improvements of 2.0% on AlpacaEval and 2.44% on HumanEval for LM & Code, and 10.4% on MBPP for Math & Code. TIES-Merging with M3 achieves an improvement of 4.01% for LM & Math on GSM8K. For LM & Code, TIES-Merging with M3 shows significant improvements of 3.11% on AlpacaEval, 25.61% on HumanEval, and 30.8% on MBPP. 2) Compared to Task Arithmetic, Average Merging and TIES-Merging tend to benefit more from M3. This is because both Average Merging and TIES-Merging use a fixed merging ratio of 1/2, whereas Task Arithmetic allows the merging ratio to vary within the range $[0.5,1.0]$. Consequently, the randomness introduced by M3 in the merging ratio has a more pronounced impact on Average Merging. This further highlights the critical role of an effective merging ratio in determining the performance of the merged model. 3) Yu et al. (2024) has indicated that the suboptimal results of merging WizardMath-13B with llama-2-13b-code-alpaca are due to llama-2-13b-code-alpaca not being well fine-tuned for code generation. In this context, the proposed M3 approach improves the pass@1 score on MBPP by 10.4% for the merged model of WizardMath-13B and llama-2-13b-code-alpaca. The improvement demonstrates that when one of the fine-tuned models to be merged is not well fine-tuned for the specific task, M3 can effectively unlock the potential of both models, maximizing the performance of the merged model. The M3 approach helps mitigate the impact of suboptimal fine-tuning on model merging performance.
4.3 Model Robustness
Out-of-distribution robustness
Figure 4: Performance of merged models (Math & Code, LM & Math, and LM & Code) using three model merging methods (Average Merging, Task Arithmetic, and TIES-Merging) on OOD datasets.
To ensure that the evaluation datasets are as representative as possible of OOD data, we select datasets with sufficiently recent release dates and ensure they cover domains that fine-tuned LLMs have not been specifically trained on. Consequently, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing. The performance of the merged LLMs is shown in Figure 4. As illustrated in Figure 4, M3 consistently enhances the performance of merged models—Math & Code, LM & Math, and LM & Code—on OOD datasets. Specifically, when Task Arithmetic is combined with M3, the Math & Code model demonstrates a 1.9% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 1.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model shows a significant 6% boost in accuracy on LiveBench-TypoFixing. Similarly, when Average Merging is combined with M3, the Math & Code model attains a 1.5% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 0.7% increase in pass@1 on LiveBench-Coding, and the LM & Code model exhibits a 4% enhancement in accuracy on LiveBench-TypoFixing. Finally, when TIES-Merging is applied alongside M3, the Math & Code model achieves a 1.1% improvement in win rate on LiveBench-Instruction, the LM & Math model records a 0.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model demonstrates a remarkable 14% improvement in accuracy on LiveBench-TypoFixing. These results underscore the robustness and versatility of M3 in enhancing model performance across diverse merging strategies and OOD tasks.
Adversarial robustness
Model Dataset Use Mixup Use Attack Metric (%) PDR (%)
Math & Code SST2 No No 57.68 38.97
Yes 35.21
Yes No 86.24 35.77
Yes 55.39
CoLA No No 45.54 98.53
Yes 0.67
Yes No 71.72 6.42
Yes 67.11
LM & Math SST2 No No 92.78 29.05
Yes 65.83
Yes No 91.28 34.55
Yes 59.75
CoLA No No 79.19 8.84
Yes 72.20
Yes No 80.54 4.52
Yes 76.89
LM & Code SST2 No No 10.55 38.04
Yes 6.54
Yes No 73.17 7.68
Yes 67.55
CoLA No No 74.21 47.42
Yes 39.02
Yes No 74.78 31.67
Yes 51.10
Table 2: Adversarial Robustness of Merged Models on the SST2 and CoLA Datasets when Executing StressTest prompt attack method. The best and second-best results are marked in bold and underlined fonts.
We employ three Prompt Attack Methods supported by the promptbench codebase (DeepWordBug, BERTAttack, and StressTest) (Zhu et al., 2024) to evaluate the adversarial robustness of three merged models (Math & Code, LM & Math, and LM & Code) obtained through the task arithmetic method. To balance experimental effectiveness with computational efficiency, we randomly selected the positions of the three attacked words in the prompts when executing the DeepWordBug and BERTAttack attacks. Adversarial robustness is assessed using the Performance Drop Rate (PDR) (Zhu et al., 2023), where a lower PDR indicates stronger robustness. Further details on PDR can be found in Appendix D. The performance of the merged LLMs is shown in Table 2.
As shown in Table 2, M3 improves the adversarial robustness of the merged models in most cases with the StressTest Prompt Attack Method. For example, with M3, the PDR of Math & Code decreased by 3.2% on the SST2 dataset and by 92.12% on the CoLA dataset, while the PDR of LM & Code decreased by 30.36% on SST2 and by 15.75% on CoLA. Furthermore, in most cases, M3 not only improves the adversarial robustness of the merged models but also enhances their performance metrics (accuracy and MCC) on the SST2 and CoLA datasets. Specifically, with M3, Math & Code demonstrates a 28.56% increase in accuracy on SST2 and a 26.18% increase in MCC on CoLA, while LM & Code achieves a 62.62% increase in accuracy on SST2. These results show that M3 effectively enhances both the adversarial robustness and the performance of the merged models in sentiment analysis and grammar correctness tasks. Detailed experimental results for the remaining Prompt Attack Methods (DeepWordBug and BERTAttack) are presented in Appendix B.
4.4 Mixup Model Merge with DARE
DARE is a model sparsification method proposed by (Yu et al., 2024), with a more detailed introduction provided in Appendix E. We combine M3 and DARE with three model merging techniques, including Average Merging, Task Arithmetic, and TIES-Merging, to compare the effects of M3 and DARE individually and explore their combined impact. The experimental results are presented in Table 1. Additionally, in the DARE method, the drop rate hyperparameter is set to 0.2.
From Table 1, we conclude that: 1) In most cases, M3 outperforms DARE, with a particularly significant advantage on certain datasets. For instance, the Math & Code model achieves a pass@1 score of 9.8% on the MBPP dataset when combined with DARE, while this score increases to 19% when combined with M3. This demonstrates that M3 unlocks new potential in model merging by randomly generating merging ratios, leading to performance improvements that surpass those achieved by DARE. 2) Combining DARE and M3 generally results in better model merging performance. For example, the LM & Math and LM & Code models, enhanced by TIES-Merging with M3 and DARE, achieve the best performance on the test datasets. While only incorporating TIES-Merging with M3 to these models, the enhanced models achieve the second best performance. This suggests that M3 and DARE can complement each other. Moreover, in some cases, M3 alone can deliver the best results, while DARE alone only achieves the best performance in very few cases, further demonstrating the superiority of M3.
5 Conclusion
Inspired by the mixup method and the Vicinal Risk Minimization (VRM) principle, we propose Mixup Model Merge (M3), a novel approach for merging fine-tuned LLMs by introducing randomness into the parameters linear interpolation process. Unlike traditional methods such as average merging and task arithmetic, M3 leverages a Beta distribution to dynamically adjust the merging ratio, enabling more flexible exploration of the parameter space. Experimental results demonstrate that M3 not only significantly enhances the performance of the merged model across various tasks but also improves its OOD and adversarial robustness. Furthermore, when combined with sparsification techniques such as DARE, our approach achieves even more favorable model merging outcomes. In summary, M3 is a simple yet powerful technique that requires minimal computational resources. By merely adjusting the merging ratio, it produces a merged model with enhanced task-specific capabilities and robustness. This exciting discovery paves the way for further research into optimizing merging ratio selection in model merging processes.
6 Limitations
There are several limitations of the M3 method: While it performs well for merging two models, (1) its scalability when merging a larger number of models, especially those with significant differences, remains uncertain. Additionally, (2) due to the inherent randomness in the merging process, multiple attempts may be required to achieve a merged model that meets expectations. This unpredictability can lead to increased computational costs, particularly in large-scale applications, resulting in a significant rise in resource consumption. Finally, (3) Our method may also be extended to a wider range of applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) models to reduce the alignment tax (FINE-TUNING, ).
Appendix A Detailed Experimental Settings
A.1 Task-Specific Fine-Tuned LLMs and Datasets Details
We conduct model merging experiments using three task-specific LLMs fine-tuned from Llama-2-13b:
  • WizardLM-13B is an instruction-following model based on Llama-2-13b, designed to improve open-domain instruction-following. Using the Evol-Instruct method (Xu et al., 2024), it generates high-complexity instruction data to reduce human annotation and enhance generalization. The model undergoes supervised fine-tuning with AI-generated data, followed by refinement via RLHF. Evaluation results show that Evol-Instruct-generated instructions outperform human-written ones, and WizardLM-13B surpasses ChatGPT in high-complexity tasks. In GPT-4 automated evaluation, it achieves over 90% of ChatGPT’s performance in 17 out of 29 tasks, demonstrating the effectiveness of AI-evolved instruction fine-tuning (Xu et al., 2024).
  • WizardMath-13B, optimized from Llama-2-13b, is designed for mathematical reasoning and enhances Chain-of-Thought (CoT) (Wei et al., 2022) capabilities. It uses Reinforcement Learning from Evol-Instruct Feedback to evolve math tasks and improve reasoning. Trained on GSM8K and MATH datasets, it excels in both basic and advanced math problems. In evaluations, WizardMath-Mistral 7B outperforms all open-source models with fewer training data, while WizardMath 70B surpasses GPT-3.5-Turbo, Claude 2, and even early GPT-4 versions in mathematical reasoning tasks.
  • llama-2-13b-code-alpaca is a code generation model fine-tuned from Llama-2-13b, designed to enhance code understanding and generation. It follows the same training approach as Stanford Alpaca (Taori et al., 2023) but focuses on code-related tasks. The model is fine-tuned with 20K instruction-following code samples generated using the Self-Instruct method (Wang et al., 2022). However, as it has not undergone safety fine-tuning, caution is required when using it in production environments.
We use one dataset to evaluate the instruction-following task:
  • AlpacaEval (Li et al., 2023) is an LLM-based automated evaluation metric that assesses model performance by testing on a fixed set of 805 instructions and computing the win rate of the evaluated model against a baseline. The evaluation process involves an LLM-based evaluator that compares the responses and determines the probability of preferring the evaluated model. In this paper, we use AlpacaEval 2.0 (Dubois et al., 2024). To reduce costs, we use chatgpt_fn for evaluation.
We use two dataset to evaluate the mathematical reasoning task:
  • GSM8K is a dataset of 8.5K high-quality, linguistically diverse grade school math word problems, designed to evaluate the multi-step mathematical reasoning abilities of large language models. It consists of 7.5K training problems and 1K test problems. In this paper, we use the 1K test set for evaluation (Cobbe et al., 2021).
  • MATH is a dataset containing 12,500 competition-level math problems, designed to evaluate and enhance the problem-solving abilities of machine learning models. It consists of 7,500 training problems and 5,000 test problems. We use the 5,000 test set for evaluation (Hendrycks et al., 2021).
We used two dataset to evaluate the code generation task:
  • HumanEval is a dataset consisting of 164 hand-written programming problems, designed to evaluate the functional correctness of code generation models. Each problem includes a function signature, docstring, function body, and unit tests. The dataset tests models’ language comprehension, reasoning, and algorithmic abilities (Chen et al., 2021).
  • MBPP is a dataset containing 974 programming problems designed to evaluate a model’s ability to synthesize Python programs from natural language descriptions. The problems range from basic numerical operations to more complex tasks involving list and string processing. The test set consists of 500 problems, which are used for evaluation in this paper (Austin et al., 2021).
A.2 Hyperparameter Setting Details in Model Merging Methods
Table 3 presents the hyperparameter search ranges for the model merging methods. For Task Arithmetic and TIES-Merging, the scaling terms are selected from $[0.5,1.0]$, while in TIES-Merging, the retain ratio for the largest-magnitude parameters is chosen from $[0.5,0.7,0.9]$. In contrast, the Average Merging method does not require any hyperparameters.
Merging Methods Search Ranges of Hyperparameters
Task Arithmetic Scaling term for merging model parameters:
[0.5, 0.6, 0.7, 0.8, 0.9, 1.0]
TIES-Merging Scaling term for merging model parameters:
[0.5, 0.6, 0.7, 0.8, 0.9, 1.0]
Ratio for retaining parameters with the
largest-magnitude values: [0.5, 0.7, 0.9]
Table 3: Hyperparameter search ranges for model merging methods.
Table 4 presents the optimal hyperparameter settings for the TIES-Merging model merging method obtained through searching. These settings are further applied to model merging experiments involving M3 and DARE.
Merging Method Model Hyperparameter Values
TIES-Merging LM & Math scaling_term=0.5, retain_ratio=0.5
LM & Code scaling_term=1.0, retain_ratio=0.7
Math & Code scaling_term=1.0, retain_ratio=0.5
Table 4: Hyperparameter settings in TIES-Merging.
A.3 Out-of-Distribution Dataset Selection Details
LiveBench (White et al., 2024) is a dynamic benchmark for large language models, featuring frequently updated questions and diverse tasks. To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench, respectively, deliberately avoiding the fine-tuning domains of the merged fine-tuned models. These tasks were released after November 2023, whereas WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca were all introduced earlier. Furthermore, their shared Llama-2-13b backbone was trained on data only up to July 2023. Consequently, these factors collectively ensure that the evaluation remains OOD in the temporal dimension.
When assessing the OOD robustness of LM & Code using the Language Comprehension category in LiveBench, only the typo-fixing task is considered. This decision is based on the fact that LiveBench is highly challenging, and the merged model performs poorly on other tasks in this category, with accuracy close to zero, rendering the evaluation results inconclusive and uninformative.
Finally, we acknowledge the limitations of these datasets. For large models like Llama-2-13b, identifying truly OOD datasets is difficult, as their training data likely covers similar distributions. These datasets are better described as "out-of-example", representing instances not explicitly seen during training. As discussed in (Wang et al., 2023), distribution shifts can occur across domains and time. While Llama-2-13b may have been trained on datasets for tasks like instruction-following, coding, and language comprehension, the datasets we selected remain valuable for OOD evaluation by capturing temporal shifts, providing insights into robustness over time.
A.4 Adversarial Robustness Evaluation Experiments Setting Details
PromptBench (Zhu et al., 2024) is a unified library designed for evaluating LLMs, providing a standardized and extensible framework. It includes several key components such as prompt construction, prompt engineering, dataset and model loading, adversarial prompt attacks, dynamic evaluation protocols, and analysis tools.
We use the Adversarial Prompt Attacks module in PromptBench aims to evaluate the robustness of LLMs against adversarial prompts. We employ three methods to perform adversarial attacks on prompts to evaluate the adversarial robustness of the merged models: DeepWordBug (Gao et al., 2018), BERTAttack (Li et al., 2020), and StressTest (Naik et al., 2018), representing Character-level, Word-level, and Sentence-level attacks, respectively.
  • DeepWordBug introduces subtle character-level perturbations (e.g., adding, deleting, or replacing characters) to words in text to deceive language models. It aims to evaluate a model’s robustness against small typographical errors that may alter the model’s performance without being easily detected.
  • BERTAttack manipulates text at the word level by replacing words with contextually similar synonyms to mislead large language models. This method tests the model’s ability to maintain accuracy despite small lexical changes that might alter the meaning of the input.
  • StressTest appends irrelevant or redundant sentences to the end of a prompt to distract and confuse language models. It assesses the model’s ability to handle extraneous information and maintain accuracy when faced with unnecessary distractions.
(a) The operational steps of TIES-Merging.
(b) After introducing M3, the Disjoint Merge step in the TIES-Merging procedure.
Figure 5: The difference between M3 and the original TIES-Merging is that, in the Disjoint Merge step, when two task vectors are retained for a given parameter, the mean of the task vectors is replaced by a random linear interpolation, while the other operations remain unchanged.
The evaluation is conducted on the Sentiment Analysis dataset (SST2 (Socher et al., 2013)) and the Grammar Correctness dataset (CoLA (Warstadt, 2019)):
  • SST2 (Socher et al., 2013): A sentiment analysis dataset designed to assess whether a given sentence conveys a positive or negative sentiment.
  • CoLA (Warstadt, 2019): A dataset for grammar correctness, where the model must determine whether a sentence is grammatically acceptable.
Appendix B Additional Experimental Results on Adversarial Robustness
Model Dataset Use Mixup Use Attack Metric (%) PDR (%)
Math & Code SST2 No No 57.68 11.73
Yes 50.92
Yes No 78.21 37.10
Yes 49.20
CoLA No No 72.87 56.97
Yes 31.35
Yes No 74.02 58.94
Yes 30.39
LM & Math SST2 No No 92.78 2.60
Yes 90.37
Yes No 91.28 3.77
Yes 87.84
CoLA No No 79.19 4.96
Yes 75.26
Yes No 80.54 1.07
Yes 79.67
LM & Code SST2 No No 10.55 98.91
Yes 0.11
Yes No 73.17 97.65
Yes 1.72
CoLA No No 74.21 8.79
Yes 67.69
Yes No 73.922 11.15
Yes 65.68
Table 5: Adversarial robustness of merged models on the SST2 and CoLA datasets when executing the DeepWordBug prompt attack method.
Model Dataset Use Mixup Use Attack Metric (%) PDR (%)
Math & Code SST2 No No 57.68 13.92
Yes 49.66
Yes No 78.21 4.11
Yes 75.00
CoLA No No 45.54 13.47
Yes 39.41
Yes No 71.72 17.25
Yes 59.35
LM & Math SST2 No No 92.78 2.22
Yes 90.71
Yes No 91.28 0.0
Yes 91.28
CoLA No No 79.19 12.00
Yes 69.70
Yes No 80.54 5.83
Yes 75.84
LM & Code SST2 No No 10.55 95.65
Yes 0.46
Yes No 73.17 55.02
Yes 32.91
CoLA No No 74.21 7.24
Yes 68.84
Yes No 73.92 7.52
Yes 68.36
Table 6: Adversarial robustness of merged models on the SST2 and CoLA datasets when executing the Bertattack prompt attack method.
All the merged models are obtained using the Task Arithmetic method. Table 5 presents the detailed experimental results of the adversarial robustness of merged models on the SST2 and CoLA datasets applying the DeepWordBug prompt attack method. Table 6 presents the detailed experimental results of the adversarial robustness of merged models on the SST2 and CoLA datasets applying the BERTAttack prompt attack method.
Appendix C Integrating M3 into the TIES-Merging Model Merging Method
Figure 5 shows the specific implementation approach to incorporating M3 into TIES-Merging. After the steps of trimming parameters with lower magnitudes and resolving sign disagreements, the two models to be merged are denoted as $M_{1}$ and $M_{2}$. During the M3 process, only the parameters that are preserved in both $M_{1}$ and $M_{2}$ are interpolated according to the model merging hyperparameter $\lambda_{m}$ to obtain the merged parameters. For parameters that are preserved in only one of the models, no interpolation is performed, and the original value from the preserved model is retained in the merged model.
Appendix D Performance Drop Rate (PDR) for Adversarial Robustness
The adversarial robustness is evaluated using the Performance Drop Rate (PDR) (Zhu et al., 2023), which is defined as follows:
$\text{PDR}=\frac{\text{Metric}_{\text{no attack}}-\text{Metric}_{\text{attack}}}{\text{Metric}_{\text{no attack}}},$ (9)
where $\text{Metric}_{\text{no attack}}$ denotes the performance metric without any prompt attack, and $\text{Metric}_{\text{attack}}$ represents the performance metric under the prompt attack. A smaller PDR indicates stronger adversarial defense against prompt attacks, implying better adversarial robustness.
Appendix E Detailed Introduction to DARE
DARE (Drop And REscale) (Yu et al., 2024) is a model sparsification method designed to reduce the redundancy of delta parameters in fine-tuned models while preserving their task-specific capabilities. In SFT, model parameters are optimized to unlock abilities for specific tasks, with the difference between fine-tuned and pre-trained parameters referred to as delta parameters.
However, studies have shown that delta parameters are often highly redundant. DARE addresses this redundancy by randomly dropping a proportion $p$ of delta parameters (referred to as the drop rate) and rescaling the remaining ones by a factor of $1/(1-p)$. This simple yet effective approach enables DARE to eliminate up to 99% of delta parameters with minimal impact on model performance, particularly in large-scale models, and it can be applied using only CPUs.
Beyond sparsification, DARE serves as a versatile plug-in for merging multiple homologous fine-tuned models (fine-tuned from the same base model) by reducing parameter interference. When combined with existing model merging techniques such as Average Merging, Task Arithmetic, and TIES-Merging, DARE facilitates the fusion of models while retaining or even enhancing task performance across multiple benchmarks. This effect is especially pronounced in decoder-based LMs, where DARE boosts task generalization.
Experiments on AlpacaEval, GSM8K, and MBPP reveal that the merged LM has the potential to outperform any individual source LM, presenting a significant new discovery. Notably, the 7B model obtained through DARE merging, SuperMario v2, ranks first among models of the same scale on the Open LLM Leaderboard (Beeching et al., 2023). These improvements were achieved without the need for retraining, positioning DARE as an efficient and resource-friendly solution for model merging.

### v2

@@
 Abstract
-Model merging integrates the parameters of multiple models into a unified model, combining their diverse capabilities.
-Existing model merging methods are often constrained by fixed parameter merging ratios.
-In this study, we propose Mixup Model Merge (M3), an innovative approach inspired by the Mixup data augmentation technique.
-This method merges the parameters of two large language models (LLMs) by randomly generating linear interpolation ratios, allowing for a more flexible and comprehensive exploration of the parameter space.
-Extensive experiments demonstrate the superiority of our proposed M3 method in merging fine-tuned LLMs: (1) it significantly improves performance across multiple tasks, (2) it enhances LLMs’ out-of-distribution (OOD) robustness and adversarial robustness, (3) it achieves superior results when combined with sparsification techniques such as DARE, and (4) it offers a simple yet efficient solution that does not require additional computational resources.
-In conclusion, M3 is a simple yet effective model merging method that significantly enhances the performance of the merged model by randomly generating contribution ratios for two fine-tuned LLMs.
+Model merging aims to integrate multiple task-specific models into a unified model that inherits the capabilities of the task-specific models, without additional training.
+Existing model merging methods often lack consideration of the varying contribution ratios of different task-specific models to the final merged model.
+In this paper, we propose Mixup Model Merge (M3), a simple yet effective method inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique.
+M3 performs randomized linear interpolation in parameter space between two task-specific LLMs, where interpolation coefficients are sampled from a Beta distribution to explore diverse contribution ratios.
+This controllable randomness allows M3 to outperform standard equal-ratio merging by discovering better contribution ratio combinations.
+Extensive experiments show that M3 significantly (1) improves merged LLM performance across tasks, (2) enhances out-of-distribution and adversarial robustness, and (3) outperforms the positive effects of the sparsification method DARE on model merging and can be further combined with DARE to achieve superior results.
+By tuning the Beta distribution’s shape parameters, (4) M3 balances exploration efficiency and diversity in contribution ratios.
 The code is available at https://github.com/MLGroupJLU/MixupModelMerge.
@@
 (b) (Left) Performance of merged models with/without M3 on OOD datasets. (Right) Adversarial robustness (PDR) of merged models with/without M3.
-(c) The performance trend of the merged models obtained through the following methods: TIES-Merging without M3 and DARE, TIES-Merging with DARE, TIES-Merging with M3, and TIES-Merging with both M3 and DARE.
-Figure 1: (a) M3 significantly boosts the model merging performance.
-(b) OOD robustness, and adversarial robustness of the merged models.
-(c) Combined with DARE, M3 delivers even better results.
-LM, MATH, and Code refer to the WizardLM13B, WizardMath-13B, and llama-2-13b-code-alpaca models.
-TA stands for Task Arithmetic and TIES stands for TIES-Merging.
+Figure 1: (a) M3 significantly boosts the model merging performance. (b) OOD robustness, and adversarial robustness of the merged models.
 Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al., 2021; Ding et al., 2023; Xia et al., 2024).
@@
 To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al., 2024; Akiba et al., 2024).
-It effectively reduces the resource-intensive demands of SFT while preserving and even enhancing model performance.
-A simple analogy for model merging is the Super Mario game, where the protagonist gains special abilities by absorbing power-up items.
-Similarly, merging model parameters integrates the strengths of different models, enabling more effective multi-task learning (Yu et al., 2024).
-However, existing model merging methods have some limitations, these methods heavily rely on predefined or heuristic parameter fusion strategies (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024) such that they fail to fully explore the parameter space, thereby restricting the potential of the merged model.
-To address this issue, we draw inspiration from Mixup (Zhang, 2017) and propose a novel technique called Mixup Model Merge (M3).
-This method introduces randomness by dynamically adjusting the contribution ratios between models, making the model merging process more flexible and enabling a thorough exploration of the parameter space.
-M3 further unlocks the potential of model merging, significantly enhancing the generalization performance of the merged model while improving its out-of-distribution (OOD) and adversarial robustness (Wang et al., 2023; Zhu et al., 2023).
+However, Existing model merging methods often adopt identical contribution ratios for different task-specific models in the merging process (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024), or rely on a few manually predefined contribution ratio values (Lin et al., 2023).
+This restricts the exploration of how varying contribution ratios influence the final merged model, and limits the ability to fully explore the resulting parameter space.
+Consequently, the potential of the merged model is significantly constrained.
+To address the above issue, we propose a novel approach, Mixup Model Merge (M3), for exploring appropriate contribution ratios among task-specific models.
+This method is inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique (Zhang, 2017).
+In contrast to conventional Mixup, M3 performs linear interpolation in the model parameter space, where the interpolation coefficients are randomly sampled from a Beta distribution.
+Specifically, this method controls the contribution ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$.
+By adjusting the parameter $\alpha$, we can precisely control the distribution of $\lambda_{m}$, enabling M3 to balance the diversity of $\lambda_{m}$ values with the efficiency of discovering suitable contribution ratios.
+In the parameter space, our method explores the family of models formed along the linear path between two task-specific LLMs by randomly sampling interpolation coefficients.
+This allows us to identify suitable interpolation coefficients (i.e., contribution ratios), thereby obtaining a merged model with superior performance.
 As shown in Figure 2, the implementation of M3 is similar to the process of proportionally mixing magical potions in Harry Potter.
-Specifically, this method dynamically controls the parameter fusion ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$.
-By adjusting the parameter $\alpha$, we can precisely control the distribution of $\lambda_{m}$, allowing M3 to explore the parameter space of model merging more deeply and thus fully unleash the potential of the models, leading to improved overall performance.
-We conducted extensive experiments with three fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively.
-Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang, 2017), we further performed comprehensive evaluations on LiveBench (White et al., 2024) and PromptBench (Zhu et al., 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness of merged models.
-The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)), and boost model merging when combined with sparsification techniques like DARE (Yu et al., 2024) (as shown in Figure 1(c)).
+We conducted extensive experiments with three homologous task-specific fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively.
+Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang, 2017), we further performed comprehensive evaluations on LiveBench (White et al., 2024) and PromptBench (Zhu et al., 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness (Wang et al., 2023; Zhu et al., 2023) of merged models.
+The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), and enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)).
 Figure 2: Implementation of M3: A process analogous to proportionally mixing magical potions in harry potter.
@@
 Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al., 2022), significantly improving the performance of various model merging techniques.
-Mixup
-Mixup is proposed to enhance the generalization ability of deep learning models by surpassing traditional Empirical Risk Minimization (ERM) (Zhang, 2017).
-It is a simple, data-agnostic augmentation technique that trains models using virtual examples created by linearly interpolating pairs of random examples and their corresponding labels.
-Rooted in the Vicinal Risk Minimization (VRM) principle (Chapelle et al., 2000), this approach improves generalization across a variety of datasets (Russakovsky et al., 2015; Krizhevsky et al., 2009; Warden, 2017; Asuncion et al., 2007), and helps reduce overfitting, sensitivity to adversarial examples, and training instability, all with minimal computational cost.
-Given two samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$, Mixup generates a new sample using the following formulas:
-[…]
-$\tilde{x}$ and $\tilde{y}$ represent the generated synthetic sample and its label, respectively, with $\lambda$ determining their interpolation ratio, typically ranging from 0 to 1.
-Here, $\lambda$ is a hyperparameter sampled from a Beta distribution, i.e., $\lambda\sim\text{Beta}(\alpha,\alpha)$, where $\alpha$ controls the strength of the interpolation between feature-target pairs.
-Inspired by Mixup, we propose a novel model merging method, M3, which generates the parameters of the merged model by performing linear interpolation between the parameters of two fine-tuned LLMs, with the interpolation ratio being random.
+DELLA (Deep et al., 2024) is a novel model merging technique that integrates a new pruning strategy called MAGPRUNE, which samples delta parameters based on their magnitudes.
+MAGPRUNE demonstrates improvements over existing methods such as DARE and TIES.
 3 Methodology
-3.1 Model Merging Problem
-Following Yu et al.
-(2024), we focus on merging fine-tuned LLMs that have been optimized from the same pre-trained backbone (Touvron et al., 2023).
-Specifically, we aim to fuse the parameters of these LLMs to create a unified model capable of handling multiple tasks.
-In this context, we restrict our attention to the merging of two models, as the mixup theory, which forms the basis of our approach, is generally applied to two entities.
-Given two tasks $t_{1}$ and $t_{2}$ with the corresponding fine-tuned LLMs having parameters $\theta_{\text{SFT}}^{t_{1}}$ and $\theta_{\text{SFT}}^{t_{2}}$, model merging aims to combine the parameters of these two models into a single model with parameters $\theta_{M}$.
-The resulting model should be able to effectively perform both tasks simultaneously, leveraging the knowledge learned from each individual model.
+We focus on merging the parameters of two homologous task-specific fine-tuned LLMs to construct a unified model that effectively preserves the capabilities of both, without requiring any further training.
 3.1 Mixup Model Merge
-Inspired by Mixup (Zhang, 2017), we propose a simple yet effective model merging method called Mixup Model Merge (M3).
-Unlike existing methods that use fixed merging ratios, M3 generates the model contribution ratio randomly, harnessing randomness to inject fresh vitality into the model merging process.
-Specifically, M3 further explores the parameter space of the merged model to unlock the potential of model merging.
-To further elaborate, given two fine-tuned LLMs with parameters $\{\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}}\}$, we combine M3 with established model merging methods to fuse these parameters and obtain a single merged model with parameters $\theta_{M}$.
+Inspired by the random interpolation strategy of Mixup (Zhang, 2017), we extend the idea from input space to parameter space, proposing Mixup Model Merge (M3).
+We consider two task-specific fine-tuned language models, denoted as $\theta_{t_{1}}^{\text{SFT}}$ and $\theta_{t_{2}}^{\text{SFT}}$, obtained by independently fine-tuning a shared pretrained base model $\theta_{\text{PRE}}$ on two different tasks $t_{1}$ and $t_{2}$, respectively.
+To explore the continuum of models between these two experts, we define a linear interpolation over their parameters:
+$\theta_{M}=\lambda_{m}\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\theta_{\text{SFT}}^{t_{2}},\quad\lambda_{m}\in(0,1)$
+This defines a one-dimensional family of interpolated models.
+Alternatively, the interpolation can be expressed in the delta space, where we defined as the parameter offsets between the task-specific fine-tuned model and the pretrained model:
+$\delta_{t_{i}}=\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}},\quad i\in\{1,2\}$
+Then the interpolated model can be reformulated as:
+$\theta_{M}=\theta_{\text{PRE}}+\lambda_{m}\delta_{t_{1}}+(1-\lambda_{m})\delta_{t_{2}}$
+In both formulations, the resulting model $\theta_{M}$ lies on the line segment connecting the two task-specific models in parameter space.
+This construction naturally defines a one-dimensional model family:
+$\mathcal{F}_{M}=\{\theta_{M}^{\lambda_{m}}\mid\lambda_{m}\in(0,1)\}$
+Such a model family allows us to study how task behaviors interact and mix at the parameter level, and serves as the foundation for our proposed Mixup Model Merge ($M^{3}$).
+Linear Interpolation Coefficient
+A key component of the M3 method is the stochastic sampling of the interpolation coefficient $\lambda_{m}$ from a Beta distribution, i.e., $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$.
+Defined over the interval $(0,1)$, the Beta distribution is naturally suited for interpolation and introduces controllable randomness into the merging process.
+Its tunable shape parameter $\alpha$ allows us to adjust the sampling bias of $\lambda_{m}$, enabling targeted and efficient exploration of the one-dimensional model space $\mathcal{F}_{M}$ for optimal model combinations.
+Specifically, as illustrated in Figure 3:
+  • When $\alpha<1$, the distribution of $\lambda_{m}$ exhibits a bimodal shape, with higher probabilities near the extremes (0 and 1), indicating that the merged model is more likely to be dominated by one of the two task-specific models.
+As $\alpha$ decreases, the range of sampled $\lambda_{m}$ values becomes more concentrated, leading to higher exploration efficiency but lower diversity of the merged models explored.
+  • When $\alpha>1$, the distribution of $\lambda_{m}$ is concentrated around the middle region (e.g., near 0.5), resulting in more balanced contributions from both models.
+As $\alpha$ increases, the range of sampled $\lambda_{m}$ values also becomes more concentrated, improving exploration efficiency while reducing the diversity of merged models.
+  • When $\alpha=1$, $\lambda_{m}$ follows a uniform distribution, meaning that all values within the interval $(0,1)$ are equally likely to be sampled.
+As $\alpha$ approaches 1, the distribution of $\lambda_{m}$ becomes more dispersed, leading to lower exploration efficiency but higher diversity in the merged models explored.
+Comparison with Grid Search
+As shown in Table 1, we compare the performance of M3 with that of grid search.
+Grid search requires manually specifying multiple fixed and discrete interpolation coefficients (e.g., $\lambda=0.1,0.2,\ldots,0.9$).
+While this approach offers fine-grained control, it lacks the ability to explore continuous values.
+Achieving finer resolution (e.g., identifying values beyond the first decimal place) requires a smaller step size, which significantly reduces search efficiency.
+In contrast, our method enables more efficient and diverse exploration of continuous interpolation coefficients.
+For example, our approach can sample values such as $\lambda_{m}=0.62156685$, which are difficult to obtain with coarse grid search.
+Empirically, we find that the model’s performance is often sensitive to the second decimal place of $\lambda$.
+To achieve such granularity, grid search would require a step size of 0.01, resulting in a prohibitively large number of evaluations.
+Our method, by comparison, typically requires only a single pass over a small set of hyperparameter values for $\alpha$ (e.g., $\alpha\in\{0.2,0.4,0.5,1,2,3,5\}$).
+In cases where a suitable $\lambda_{m}$ is not found in the first pass, we perform a few additional samplings under a selected $\alpha$ value—no more than three times in our experiments.
+In Table 2, we report the specific $\lambda_{m}$ values that lead to performance improvements, along with the corresponding distribution shape parameter $\alpha$ and the total number of sampling iterations.
+Method Distribution of $\lambda_{m}$ Continuity of $\lambda_{m}$ Efficiency Diversity Controllability
+Grid Search Fixed discrete values (e.g., 0.1–0.9) No Low Medium High
+M3+Beta(1,1) Uniform over (0,1) Yes Medium High Medium
+M3+Beta(5,5) Concentrated near the center (0.5) Yes High Low High
+M3+Beta(0.2,0.2) Concentrated near the endpoints (0 or 1) Yes High Low High
+Table 1: Comparison of interpolation coefficient $\lambda_{m}$ sampling strategies.
+M3 Implementation
+M3 is a plug-and-play approach that can be integrated into existing representative model merging methods to enhance the performance of the resulting merged models.
 As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022).
@@
 $=\theta_{\text{PRE}}+\lambda\cdot\sum_{i=1}^{2}(\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}}),$ (2)
-where $\theta_{\text{PRE}}\in\mathbb{R}^{d}$ represents the parameters of the pre-trained language model (PLM), such as Llama 2 (Touvron et al., 2023).
-$\lambda$ is a scaling factor that weights the contribution of each model during the merging process.
-$\delta^{t_{i}}$ denotes the delta parameter (Ilharco et al., 2022), which is defined as the difference between the parameters of the language models (LMs) before and after SFT, i.e., $\delta^{t}=\theta_{\text{SFT}}^{t}-\theta_{\text{PRE}}\in\mathbb{R}^{d}$, where $t$ refers to task $t$.
+$\lambda$ is a scaling term that takes values from the set $\{0.0,0.1,\ldots,1.0\}$.
 When introducing M3, the process for Average Merging is reformulated as:
@@
 $\theta_{M}$ $=\theta_{\text{PRE}}+\lambda_{m}\delta^{t_{1}}+(1-\lambda_{m})\delta^{t_{2}},$ (4)
-where $\lambda_{m}$ determines the linear interpolation ratio between the two fine-tuned LLMs, and is generally a value between 0 and 1.
-$\lambda_{m}$ is sampled from a Beta distribution, typically $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$, where $\alpha$ controls the shape of the Beta distribution.
-The hyperparameter $\alpha$ for M3 is selected from the range $[0.2,0.4,0.5,1,2,3,5]$.
-As shown in Figure 3: (1) When $\alpha=1$, $\lambda$ follows a uniform distribution, meaning all values within the range $(0,1)$ are equally likely to be sampled.
-(2) When $\alpha<1$, the distribution of $\lambda$ exhibits a bimodal shape, with higher probabilities near the extremes (0 and 1).
-This indicates that the merged model is more likely to be dominated by one of the two models.
-(3) When $\alpha>1$, the distribution of $\lambda$ becomes concentrated around the middle (e.g., 0.5), resulting in more balanced contributions from both models.
+where
+$\lambda_{m}\in[0,1],\quad\lambda_{m}\sim\mathrm{Beta}(\alpha,\alpha),$
+$\lambda_{m}$ controls the interpolation ratio, i.e., the contribution ratio, between the two fine-tuned models. $\alpha$ determines the shape of the Beta distribution.
 Figure 3: The Beta distribution visualization for different $\alpha$ values.
 3.2 Theoretical Analysis
-Mixup performs linear interpolations between data samples and their labels in the data space (Zhang, 2017).
-Similarly, M3 can be viewed as applying random linear interpolation in the parameter space, where the interpolation occurs between the parameters of two fine-tuned models trained on different tasks.
-M3 represents a natural extension of the Vicinal Risk Minimization (VRM) principle (Chapelle et al., 2000) into the model parameter space.
-In data space, VRM introduces a vicinal distribution to simulate the true data distribution, thereby increasing the diversity of training data and enhancing the model’s generalization ability.
-Similarly, M3 extends this principle by constructing a virtual neighborhood in the model parameter space between two task-specific models, combining their knowledge in a way that is more natural and balanced.
-In the context of model merging, interpolating between two sets of model parameters, $\theta_{\text{SFT}}^{t_{1}}$and $\theta_{\text{SFT}}^{t_{2}}$, defines a new neighborhood distribution in the parameter space, which can be expressed as:
-[…]
-where $P_{\nu}(\theta_{M})$ represents the probability distribution of the new model parameters $\theta_{M}$ generated through interpolation.
-The function $\nu(\theta_{M}\mid\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}})$ defines the range and behavior of $\theta_{M}$, depending on the interpolation strategy (such as linear interpolation).
-Additionally, $P(\theta_{\text{SFT}}^{t_{1}})$ and $P(\theta_{\text{SFT}}^{t_{2}})$ represent the prior distributions of the parameters of the two models, respectively.
-Under the linear interpolation strategy, the function $\nu(\theta_{M}\mid\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}})$ is defined as:
-$\theta_{M}=\lambda_{m}\cdot\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\cdot\theta_{\text{SFT}}^{t_{2}},$ (7)
-where $\lambda_{m}\in(0,1)$ is the interpolation ratio. The corresponding distribution $P_{\nu}(\theta_{M})$ is expressed as:
-where $\delta(\cdot)$ is the Dirac delta function, ensuring that $\theta_{M}$ satisfies the linear interpolation rule.
-By using this linear interpolation method, the parameters $\theta_{\text{SFT}}^{t_{1}}$ and $\theta_{\text{SFT}}^{t_{2}}$ are combined in the parameter space, forming a new neighborhood distribution $P_{\nu}(\theta_{M})$.
-This process effectively merges the knowledge of both models into a unified parameter space, thus achieving robust performance across different tasks while maintaining the original strengths of the models.
-By interpolating the parameters, M3 encourages the model to learn smoother decision boundaries.
-In this context, smoother decision boundaries can be understood as the boundaries between different tasks, such as instruction following, mathematical reasoning, and code generation, where the model must understand and adapt to each task differently.
-In tasks $t_{1}$ and $t_{2}$, M3 creates a virtual neighborhood that seamlessly integrates knowledge from both tasks.
-This approach prevents the merged model from overfitting task-specific details, ensuring a balanced and effective performance across all tasks.
-Secondly, M3 introduces a linear inductive bias in the parameter space, encouraging the merged model parameters to lie on a linear manifold between the two source models.
-This linear structure offers significant advantages in terms of simplicity and generalization.
-According to Occam’s Razor, simpler solutions tend to generalize better.
-By performing linear interpolation between two sets of parameters, M3 avoids unnecessary complexity in the model merging process, leading to a more straightforward and efficient solution.
-Thirdly, M3 can improve the performance of the merged model across multiple tasks by mitigating task conflicts.
-Different tasks may require conflicting parameter values, which can lead to performance degradation on one task while optimizing for another.
-Linear interpolation helps balance these conflicts, resulting in a model that performs well among all tasks.
+According to the theoretical and empirical findings of Neyshabur et al.
+(2020), fine-tuning models from the same pretrained initialization can guide them into a shared, flat, and connected low-loss basin.
+As a result, the linear interpolation path between such fine-tuned models remains within this low-loss basin and does not encounter significant performance barriers.
+In some cases, models along the interpolation path even outperform the endpoints.
+Since both $\theta_{t1}^{\mathrm{SFT}}$ and $\theta_{t2}^{\mathrm{SFT}}$ are fine-tuned from the same pretrained model, they can be considered to lie within the same flat and connected low-loss basin.
+Therefore, linearly interpolating between them is unlikely to cause a collapse in performance, and may even yield interpolated models with better performance than either endpoint.
+This provides a theoretical justification for exploring interpolated models along the linear path between two fine-tuned models as a means of finding superior merged models.
+Consider the $i$-th parameter of the interpolated model:
+$\theta_{M,i}=\theta_{\mathrm{PRE},i}+\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i},$ (5)
+where $\delta_{t1,i}=\theta_{t1,i}^{\mathrm{SFT}}-\theta_{\mathrm{PRE},i}$ and $\delta_{t2,i}=\theta_{t2,i}^{\mathrm{SFT}}-\theta_{\mathrm{PRE},i}$ denote the delta parameters from the pretrained model $\theta_{\mathrm{PRE}}$ to the task-specific fine-tuned models $\theta_{t1}^{\mathrm{SFT}}$ and $\theta_{t2}^{\mathrm{SFT}}$, respectively.
+Define a function representing the change of the merged model’s $i$-th parameter relative to the pretrained model:
+$f_{i}(\lambda_{m})=\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i}.$ (6)
+If $\delta_{t1,i}>0$ and $\delta_{t2,i}<0$, the delta parameters are in conflict and may partially cancel each other. In particular, when
+$\lambda_{m}=\frac{|\delta_{t2,i}|}{|\delta_{t1,i}|+|\delta_{t2,i}|}\in(0,1),$ (7)
+we have $f_{i}(\lambda_{m})=0$, meaning the conflicting delta parameters are exactly canceled out at that parameter.
+Even when $\lambda_{m}$ is close to this cancellation point, the magnitude of the update is significantly reduced.
+This shows that linear interpolation can effectively suppress conflicting delta parameters at certain parameters, providing stability in the merged model.
 4 Experiments
@@
 More details of these LLMs and datasets can be found in Appendix A.1.
-The Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness
-To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al., 2024), respectively.
+Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness
+To evaluate OOD robustness, we evaluate merged models including math & code, LM & math, and LM & code on instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al., 2024), respectively.
 More details on OOD benchmarks are given in Appendix A.3.
-We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al., 2024) to assess the robustness of LLMs against adversarial prompts.
+We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al., 2024) to evaluate the robustness of LLMs against adversarial prompts.
 Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al., 2018), BERTAttack (word-level) (Li et al., 2020), and StressTest (sentence-level) (Naik et al., 2018).
@@
 Implementation Details
-Unless otherwise specified, the details of the model merging experiments are consistent with Yu et al.
+In each distinct merging experiment—where “distinct” refers to variations in either the task-specific LLMs or the model merging methods—we perform a full sweep over the sole hyperparameter of M3, $\alpha$, using the set $\{0.2,0.4,0.5,1,2,3,5\}$.
+For each $\alpha$ value, we sample $\lambda_{m}$ once from the corresponding Beta distribution, resulting in a total of seven samples in each merging experiment.
+Seven samples are not the minimum required to observe significant performance gains.
+In most cases, a smaller number of samples is sufficient to identify a performant interpolation coefficient.
+However, to enable a more comprehensive exploration, we fix the number of samples to seven—each corresponding to a representative shape of the Beta distribution.
+While the stochastic nature of our method means there may be cases where none of the seven samples yields a significant improvement, such cases are not often observed in our experiments and thus not considered.
+Given the strong performance observed with just seven samples, and to conserve computational resources, we refrain from further sampling.
+Unless otherwise specified, the remaining details of the model merging experiments follow Yu et al.
 (2024).
-The hyperparameter $\alpha$ for M3 is chosen from the range $[0.2,0.4,0.5,1,2,3,5]$.
-For a detailed description of the hyperparameter settings in model merging methods, please refer to Appendix A.2.
-Additionally, all experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
+For a detailed description of the hyperparameter settings used in the representative existing model merging methods, please refer to Appendix A.2.
+All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
 4.2 Merging Task-Specific Fine-Tuned LLMs
-Merging Methods Models Use Model Mixup Use DARE Instruction Following Mathematical Reasoning Code Generating
-AlpacaEval GSM8K MATH HumanEval MBPP
-/ LM No No 45.14 2.20 0.04 36.59 34.00
-Math No No / 64.22 14.02 / /
-Code No No / / / 23.78 27.60
-Average Merging LM & Math No No 45.28 66.34 13.40 / /
-Yes No 44.40 66.26 13.80 / /
-No Yes 44.22 66.57 12.96 / /
-Yes Yes 43.53 66.57 14.12 / /
-LM & Code No No 36.60 / / 29.88 32.00
-Yes No 43.91 / / 37.20 34.40
-No Yes 38.81 / / 31.71 32.40
-Yes Yes 40.31 / / 36.59 37.00
-Math & Code No No / 56.17 10.28 8.53 8.20
-Yes No / 63.61 14.02 8.54 19.20
-No Yes / 56.18 10.28 6.10 7.80
-Yes Yes / 64.97 13.54 9.76 21.20
-Task Arithmetic LM & Math No No 45.78 66.34 13.40 / /
-Yes No 41.65 66.34 13.74 / /
-No Yes 49.00 66.64 13.02 / /
-Yes Yes 44.90 67.32 13.74 / /
-LM & Code No No 44.64 / / 32.93 33.60
-Yes No 46.64 / / 35.37 33.80
-No Yes 41.47 / / 35.98 33.00
-Yes Yes 45.20 / / 35.98 35.20
-Math & Code No No / 64.67 13.98 8.54 8.60
-Yes No / 63.53 13.94 7.93 19.00
-No Yes / 65.05 13.96 10.37 9.80
-Yes Yes / 65.13 14.32 8.54 18.00
-TIES- Merging LM & Math No No 38.63 14.56 2.12 / /
-Yes No 38.73 18.57 2.48 / /
-No Yes 37.92 18.04 2.34 / /
-Yes Yes 39.93 19.26 2.82 / /
-LM & Code No No 41.85 / / 0.0 0.0
-Yes No 44.96 / / 25.61 30.80
-No Yes 43.13 / / 0.0 0.0
-Yes Yes 45.65 / / 26.83 33.20
-Math & Code No No / 64.67 13.68 9.15 22.60
-Yes No / 64.75 14.16 9.76 21.4
-No Yes / 64.82 13.88 10.37 23.60
-Yes Yes / 64.75 14.78 9.15 19.60
-Table 1: Performance of merging task-specific LLMs WizardLM-13B (LM), WizardMath-13B (Math), and llama-2-13b-codealpaca (Code) on all the datasets.
-The best and second-best results are marked in bold and underlined fonts.
-We integrate M3 into three prominent model merging techniques: Average Merging, Task Arithmetic, and TIES-Merging. The performance of merging task-specific fine-tuned LLMs is presented in Table 1.
-From Table 1, we obtain the following observations: 1) M3 generally enhances Average Merging, Task Arithmetic, and TIES-Merging when merging fine-tuned LLMs.
+Merging Methods Models Use M3 (Ours) Sampling Details Instruction Following Mathematical Reasoning Code Generating Avg.
+$\lambda_{m}$ $\alpha$ AlpacaEval GSM8K MATH HumanEval MBPP
+- LM - - - 45.14 2.20 0.04 36.59 34.00 -
+Math - - - - 64.22 14.02 - - -
+Code - - - - - - 23.78 27.60 -
+Average Merging LM & Math No - - 45.28 66.34 13.40 - - 41.67
+Yes 0.35 2 44.40 66.26 13.80 - - 41.48
+LM & Code No - - 36.60 - - 29.88 32.00 32.82
+Yes 0.84 2 43.91$\uparrow$ - - 37.20$\uparrow$ 34.40 38.50$\uparrow$
+Math & Code No - - - 56.17 10.28 8.53 8.20 20.80
+Yes 0.92 0.4 - 63.61$\uparrow$ 14.02$\uparrow$ 8.54 19.20$\uparrow$ 26.34$\uparrow$
+Task Arithmetic LM & Math No - - 45.78 66.34 13.40 - - 41.84
+Yes 0.34 2 41.65 66.34 13.74 - - 40.58
+LM & Code No - - 44.64 - - 32.93 33.60 37.06
+Yes 0.99 0.4 46.64 - - 35.37 33.80 38.60
+Math & Code No - - - 64.67 13.98 8.54 8.60 23.95
+Yes 0.98 0.5 - 63.53 13.94 7.93 19.00$\uparrow$ 26.1
+TIES- Merging LM & Math No - - 38.63 14.56 2.12 - - 18.44
+Yes 0.62 1 38.73 18.57$\uparrow$ 2.48 - - 19.92
+LM & Code No - - 41.85 - - 0.0 0.0 13.95
+Yes 0.84 0.5 44.96 - - 25.61$\uparrow$ 30.80$\uparrow$ 33.79$\uparrow$
+Math & Code No - - - 64.67 13.68 9.15 22.60 27.57
+Yes 0.63 3 - 64.75 14.16 9.76 21.4 27.52
+DARE LM & Math No - - 49.00 66.64 13.2 - - 42.95
+Yes 0.38 0.4 44.90 67.32 13.74 - - 41.98
+LM & Code No - - 41.47 - - 35.98 33.00 36.82
+Yes 0.99 0.5 45.20$\uparrow$ - - 35.98 35.20 38.79
+Math & Code No - - - 65.05 10.37 9.15 9.80 23.59
+Yes 0.98 0.5 - 65.13 14.32$\uparrow$ 8.54 18.00$\uparrow$ 26.50$\uparrow$
+Table 2: Comparison of Merged Model Performance Before and After Applying M3.
+We use three task-specific LLMs: WizardLM-13B (LM), WizardMath-13B (Math), and llama-2-13b-codealpaca (Code).
+Bold indicates the best results for each dataset, underline highlights improvements with M3, and $\uparrow$ denotes significant gains after introducing M3.
+Our proposed method, M3, is a plug-and-play approach.
+To evaluate its effectiveness in improving merged model performance, we integrate M3 into several representative existing model merging methods (Average Merging, Task Arithmetic, TIES-Merging, DARE) and compare the performance of the merged models before and after applying M3.
+The results is presented in Table 2.
+Performance Gains of M3 Across Merging Methods and Tasks
+As shown in Table 2, M3 generally enhances existing model merging methods, achieving over 10-point improvements on multiple datasets and up to 30 points on one.
 For example, the improvements achieved by Average Merging with M3 for Math & Code are 7.43% on GSM8K, 3.74% on Math, and 11.0% on MBPP.
@@
 For LM & Code, TIES-Merging with M3 shows significant improvements of 3.11% on AlpacaEval, 25.61% on HumanEval, and 30.8% on MBPP.
-2) Compared to Task Arithmetic, Average Merging and TIES-Merging tend to benefit more from M3.
-This is because both Average Merging and TIES-Merging use a fixed merging ratio of 1/2, whereas Task Arithmetic allows the merging ratio to vary within the range $[0.5,1.0]$.
-Consequently, the randomness introduced by M3 in the merging ratio has a more pronounced impact on Average Merging.
-This further highlights the critical role of an effective merging ratio in determining the performance of the merged model.
-3) Yu et al.
-(2024) has indicated that the suboptimal results of merging WizardMath-13B with llama-2-13b-code-alpaca are due to llama-2-13b-code-alpaca not being well fine-tuned for code generation.
-In this context, the proposed M3 approach improves the pass@1 score on MBPP by 10.4% for the merged model of WizardMath-13B and llama-2-13b-code-alpaca.
-The improvement demonstrates that when one of the fine-tuned models to be merged is not well fine-tuned for the specific task, M3 can effectively unlock the potential of both models, maximizing the performance of the merged model.
-The M3 approach helps mitigate the impact of suboptimal fine-tuning on model merging performance.
+Furthermore, as evidenced by the Avg.
+column in Table 2, M3 consistently improves the overall performance of the merged models.
+This enhancement is particularly crucial for scenarios where the merged models must strike a balance between task-specific capabilities.
+A few datasets do not show performance improvements with M3, but the results remain on par with those without M3, indicating that it does not degrade overall performance.
+Moreover, such cases are negligible compared to the overall benefits brought by M3.
+As shown in Table 2, M3 achieves the best performance on four out of five task-specific datasets, surpassing even the individually fine-tuned LLMs (i.e., LM, Math, and Code).
+This suggests that M3 enables the discovery of interpolated models along linear paths between two models that outperform both endpoints, providing support for the theoretical insights discussed in Section 3.2.
+In addition, we also compare our proposed method M3 with existing model merging methods. The details are in Appendix G.
+Mitigating Suboptimal Fine-Tuning with M3
+Yu et al.
+(2024) noted that the poor performance of merging WizardMath-13B with llama-2-13b-code-alpaca stems from the latter not being sufficiently fine-tuned for code generation tasks.
+As shown in Table 2, the proposed M3 method improves the pass@1 score on the code generation dataset MBPP by 10.4%.
+This improvement demonstrates that even when one of the fine-tuned models is suboptimal for a specific task, M3 can effectively unlock the merging potential by adaptively exploring suitable interpolation coefficients (i.e., contribution ratios), thereby maximizing the performance of the merged model.
+In other words, M3 significantly mitigates the negative impact of insufficient fine-tuning on merging quality, further highlighting the critical importance of task-specific LLMs’ contribution ratio in determining the performance of the merged model.
 4.3 Model Robustness
@@
 Consequently, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing.
-The performance of the merged LLMs is shown in Figure 4.
 As illustrated in Figure 4, M3 consistently enhances the performance of merged models—Math & Code, LM & Math, and LM & Code—on OOD datasets.
@@
 Finally, when TIES-Merging is applied alongside M3, the Math & Code model achieves a 1.1% improvement in win rate on LiveBench-Instruction, the LM & Math model records a 0.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model demonstrates a remarkable 14% improvement in accuracy on LiveBench-TypoFixing.
-These results underscore the robustness and versatility of M3 in enhancing model performance across diverse merging strategies and OOD tasks.
 Adversarial robustness
@@
 Further details on PDR can be found in Appendix D.
-The performance of the merged LLMs is shown in Table 2.
 As shown in Table 3, M3 improves the adversarial robustness of the merged models in most cases with the StressTest Prompt Attack Method.
@@
 Detailed experimental results for the remaining Prompt Attack Methods (DeepWordBug and BERTAttack) are presented in Appendix B.
-4.4 Mixup Model Merge with DARE
+4.4 Comparison of Mixup Model Merge and DARE
 DARE is a model sparsification method proposed by (Yu et al., 2024), with a more detailed introduction provided in Appendix E.
-We combine M3 and DARE with three model merging techniques, including Average Merging, Task Arithmetic, and TIES-Merging, to compare the effects of M3 and DARE individually and explore their combined impact.
-The experimental results are presented in Table 1.
+This section takes DARE as an example to discuss the relationship between the existing model sparsification method and M3.
+We integrated M3 and DARE with three existing model merging techniques—Average Merging, Task Arithmetic, and TIES-Merging—to compare these two plug-and-play methods.
+Detailed experimental results are provided in Appendix F.
 Additionally, in the DARE method, the drop rate hyperparameter is set to 0.2.
-From Table 1, we conclude that: 1) In most cases, M3 outperforms DARE, with a particularly significant advantage on certain datasets.
+From Table 8, it can be concluded that M3 generally outperforms DARE, exhibiting significantly superior performance on certain datasets.
 For instance, the Math & Code model achieves a pass@1 score of 9.8% on the MBPP dataset when combined with DARE, while this score increases to 19% when combined with M3.
-This demonstrates that M3 unlocks new potential in model merging by randomly generating merging ratios, leading to performance improvements that surpass those achieved by DARE.
-2) Combining DARE and M3 generally results in better model merging performance.
-For example, the LM & Math and LM & Code models, enhanced by TIES-Merging with M3 and DARE, achieve the best performance on the test datasets.
-While only incorporating TIES-Merging with M3 to these models, the enhanced models achieve the second best performance.
+As shown in Table 8, Combining DARE and M3 generally results in better model merging performance.
 This suggests that M3 and DARE can complement each other.
@@
 5 Conclusion
-Inspired by the mixup method and the Vicinal Risk Minimization (VRM) principle, we propose Mixup Model Merge (M3), a novel approach for merging fine-tuned LLMs by introducing randomness into the parameters linear interpolation process.
-Unlike traditional methods such as average merging and task arithmetic, M3 leverages a Beta distribution to dynamically adjust the merging ratio, enabling more flexible exploration of the parameter space.
-Experimental results demonstrate that M3 not only significantly enhances the performance of the merged model across various tasks but also improves its OOD and adversarial robustness.
-Furthermore, when combined with sparsification techniques such as DARE, our approach achieves even more favorable model merging outcomes.
-In summary, M3 is a simple yet powerful technique that requires minimal computational resources.
-By merely adjusting the merging ratio, it produces a merged model with enhanced task-specific capabilities and robustness.
-This exciting discovery paves the way for further research into optimizing merging ratio selection in model merging processes.
+Addressing the issue that existing methods overlook the varying contribution ratios of different models, we propose Mixup Model Merge (M3), a novel approach based on the randomized linear interpolation strategy of Mixup.
+M3 performs linear interpolation in the parameter space between two task-specific large language models, with interpolation coefficients sampled from a Beta distribution to explore a diverse range of contribution ratios.
+This method leverages controllable randomness to significantly outperform traditional equal-ratio merging.
+Extensive experiments demonstrate that M3 not only improves merged model performance across multiple tasks but also enhances out-of-distribution and adversarial robustness, and effectively combines with model sparsification methods such as DARE.
+By tuning the Beta distribution parameters, M3 balances efficiency and diversity in exploring contribution ratios.
 6 Limitations
-There are several limitations of the M3 method: While it performs well for merging two models, (1) its scalability when merging a larger number of models, especially those with significant differences, remains uncertain.
-Additionally, (2) due to the inherent randomness in the merging process, multiple attempts may be required to achieve a merged model that meets expectations.
-This unpredictability can lead to increased computational costs, particularly in large-scale applications, resulting in a significant rise in resource consumption.
-Finally, (3) Our method may also be extended to a wider range of applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) models to reduce the alignment tax (FINE-TUNING, ).
+Despite its advantages, the M3 method has several limitations: (1) Our method is currently limited to pairwise model merging, as random linear interpolation is typically defined between two entities.
+Extending it to merge three or more models requires new strategies for sampling multi-way interpolation coefficients, which introduces additional complexity and remains an open direction for future work.
+(2) This study focuses on merging homogeneous models fine-tuned from the same pretrained backbone.
+Applying M3 to heterogeneous models with different architectures or fine-tuning paths poses new challenges and will be explored in future work.
+(3) Our method may also be extended to broader applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) models to reduce the alignment tax (FINE-TUNING, ).
 Appendix A Detailed Experimental Settings
@@
 These improvements were achieved without the need for retraining, positioning DARE as an efficient and resource-friendly solution for model merging.
+Appendix F Detailed M3 and DARE Comparison Results
+Merging Methods Models Use M3 (Ours) Use DARE Instruction Following Mathematical Reasoning Code Generating
+AlpacaEval GSM8K MATH HumanEval MBPP
+Average Merging LM & Math Yes No 44.40 66.26 13.80 - -
+No Yes 44.22 66.57 12.96 - -
+Yes Yes 43.53 66.57 14.12 - -
+LM & Code Yes No 43.91 - - 37.20 34.40
+No Yes 38.81 - - 31.71 32.40
+Yes Yes 40.31 - - 36.59 37.00
+Math & Code Yes No - 63.61 14.02 8.54 19.20$\uparrow$
+No Yes - 56.18 10.28 6.10 7.80
+Yes Yes - 64.97 13.54 9.76 21.20
+Task Arithmetic LM & Math Yes No 41.65 66.34 13.74 - -
+No Yes 49.00 66.64 13.02 - -
+Yes Yes 44.90 67.32 13.74 - -
+LM & Code Yes No 46.64 - - 35.37 33.80
+No Yes 41.47 - - 35.98 33.00
+Yes Yes 45.20 - - 35.98 35.20
+Math & Code Yes No - 63.53 13.94 7.93 19.00
+No Yes - 65.05 13.96 10.37 9.80
+Yes Yes - 65.13 14.32 8.54 18.00
+TIES- Merging LM & Math Yes No 38.73 18.57 2.48 - -
+No Yes 37.92 18.04 2.34 - -
+Yes Yes 39.93 19.26 2.82 - -
+LM & Code Yes No 44.96 - - 25.61 30.80
+No Yes 43.13 - - 0.0 0.0
+Yes Yes 45.65 - - 26.83 33.20
+Math & Code Yes No - 64.75 14.16 9.76 21.4
+No Yes - 64.82 13.88 10.37 23.60
+Yes Yes - 64.75 14.78 9.15 19.60
+Table 8: Comparison of our method M3 and DARE. The best and second-best results are marked in bold and underlined fonts.
+We compare the two plug-and-play methods, M3 and DARE, by integrating each with three existing model merging techniques—Average Merging, Task Arithmetic, and TIES-Merging.
+The detailed experimental results are presented in Table 8.
+Appendix G Comparison of M3 and Existing Model Merging Approaches
+Models Merging Methods Instruction Following Mathematical Reasoning Code Generating Avg.
+AlpacaEval GSM8K MBPP
+LM & Math Average Merging 45.28 66.34 - 55.81
+Task Arithmetic 45.78 66.34 - 56.06
+TIES-Merging 38.63 14.56 - 26.60
+DARE 49.00 66.64 - 57.82
+DELLA 44.03 48.45 - 46.24
+M3 (Ours) 44.90 67.32 - 56.11
+LM & Code Average Merging 36.60 - 32.00 34.3
+Task Arithmetic 44.64 - 33.60 39.12
+TIES-Merging 41.85 - 0.0 20.93
+DARE 41.47 - 33.00 37.24
+DELLA 40.99 - 35.00 38.00
+M3 (Ours) 45.20 - 35.20 40.2
+Math & Code Average Merging - 56.17 8.20 32.19
+Task Arithmetic - 64.67 8.60 36.64
+TIES-Merging - 64.67 22.60 43.64
+DARE - 65.05 9.80 37.43
+DELLA - 63.08 20.2 41.64
+M3 (Ours) - 64.75 21.4 43.10
+Table 9: Comparison of our method M3 and existing model merging approaches. The best and second-best results are marked in bold and underlined fonts.
+For each baseline method—Average Merging, Task Arithmetic, TIES-Merging, and DARE—we combine it with our proposed M3 method and select the best-performing combination for each merged model (LM & Math, LM & Code and Math & Code).
+These optimal results are reported as M3 (Ours) in Table 9.
+As presented in Table 9, our proposed method M3 achieves the best overall performance across the three evaluated tasks.
+Importantly, M3 attains these improvements without involving complex architectural designs or intensive computations; it merely adjusts the contribution ratios of constituent models to realize substantial gains, thereby outperforming the DELLA approach.
+DELLA advances model sparsification techniques by building upon TIES-Merging and DARE, introducing a novel pruning strategy termed MAGPRUNE, which selects delta parameters based on their magnitudes.
+This strategy has inspired us to investigate new directions for enhancing the performance of merged models.

### v3

@@
 By tuning the Beta distribution’s shape parameters, (4) M3 balances exploration efficiency and diversity in contribution ratios.
-The code is available at https://github.com/MLGroupJLU/MixupModelMerge.
-Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation
-Yue Zhou1  Yi Chang1,2,3  Yuan Wu1111Corresponding authors 1School of Artificial Intelligence, Jilin University 2Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 3International Center of Future Science, Jilin University yuezhou24@mails.jlu.edu.cn, yichang@jlu.edu.cn, yuanwu@jlu.edu.cn
+The code is provided in the supplementary materials.
+Codehttps://github.com/MLGroupJLU/MixupModelMerge
 1 Introduction
-In the field of Natural Language Processing (NLP), the emergence of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; OpenAI, 2023; Chowdhery et al., 2023) represents a revolutionary breakthrough.
-With their remarkable capabilities, these models have demonstrated outstanding performance across various tasks (Jiao et al., 2023; Chang et al., 2024b; Nam et al., 2024; Xing, 2024; Guo et al., 2024), significantly advancing NLP technologies.
-(a) Performance of the merged models obtained through Task Arithmetic and TIES-Merging with/without M3.
-(b) (Left) Performance of merged models with/without M3 on OOD datasets. (Right) Adversarial robustness (PDR) of merged models with/without M3.
-Figure 1: (a) M3 significantly boosts the model merging performance. (b) OOD robustness, and adversarial robustness of the merged models.
-Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al., 2021; Ding et al., 2023; Xia et al., 2024).
-However, SFT requires substantial computational resources and long training times (Brown et al., 2020; Chang et al., 2024a).
-To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al., 2024; Akiba et al., 2024).
-However, Existing model merging methods often adopt identical contribution ratios for different task-specific models in the merging process (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024), or rely on a few manually predefined contribution ratio values (Lin et al., 2023).
+In the field of Natural Language Processing (NLP), the emergence of large language models (LLMs) (Brown et al.
+2020; Touvron et al.
+2023; OpenAI 2023; Chowdhery et al.
+2023) represents a revolutionary breakthrough.
+With their remarkable capabilities, these models have demonstrated outstanding performance across various tasks (Jiao et al.
+2023; Chang et al.
+2024; Nam et al.
+2024; Xing 2024; Guo et al.
+2024), significantly advancing NLP technologies.
+(a) Performance of the merged models obtained through Task Arithmetic and TIES-Merging with or without M3
+(b) (Left) Performance on OOD datasets. (Right) Adversarial robustness (PDR) of merged models with or without M3
+Figure 1: (a) M3 significantly improves the performance of model merging. (b) OOD and adversarial robustness of merged models with or without M3
+Figure 2: Implementation of M3: A process analogous to proportionally mixing magical potions in Harry Potter.
+The proposed method controls the contribution ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$.
+The distribution of $\lambda_{m}$ is controlled by adjusting $\alpha$.
+Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al.
+2021; Ding et al.
+2023; Xia et al.
+2024).
+However, SFT requires substantial computational resources and long training times (Brown et al.
+2020; Chang, Chang, and Wu 2024).
+To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al.
+2024; Akiba et al.
+2024).
+However, Existing model merging methods often adopt identical contribution ratios for different task-specific models in the merging process (Wortsman et al.
+2022; Ilharco et al.
+2022; Matena and Raffel 2022; Jin et al.
+2022; Yadav et al.
+2023; Yu et al.
+2024), or rely on a few manually predefined contribution ratio values (Lin et al.
+2023).
 This restricts the exploration of how varying contribution ratios influence the final merged model, and limits the ability to fully explore the resulting parameter space.
@@
 To address the above issue, we propose a novel approach, Mixup Model Merge (M3), for exploring appropriate contribution ratios among task-specific models.
-This method is inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique (Zhang, 2017).
+This method is inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique (Zhang 2017).
 In contrast to conventional Mixup, M3 performs linear interpolation in the model parameter space, where the interpolation coefficients are randomly sampled from a Beta distribution.
@@
 This allows us to identify suitable interpolation coefficients (i.e., contribution ratios), thereby obtaining a merged model with superior performance.
-As shown in Figure 2, the implementation of M3 is similar to the process of proportionally mixing magical potions in Harry Potter.
-We conducted extensive experiments with three homologous task-specific fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively.
-Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang, 2017), we further performed comprehensive evaluations on LiveBench (White et al., 2024) and PromptBench (Zhu et al., 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness (Wang et al., 2023; Zhu et al., 2023) of merged models.
+We conducted extensive experiments with three homologous task-specific fine-tuned LLMs: WizardLM-13B (Xu et al.
+2024), WizardMath-13B (Luo et al.
+2023), and llama-2-13b-code-alpaca (Chaudhary 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively.
+Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang 2017), we further performed comprehensive evaluations on LiveBench (White et al.
+2024) and PromptBench (Zhu et al.
+2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness (Wang et al.
+2023; Zhu et al.
+2023) of merged models.
 The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), and enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)).
-Figure 2: Implementation of M3: A process analogous to proportionally mixing magical potions in harry potter.
-The proposed method controls the contribution ratio between two fine-tuned llms by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$.
-The distribution of $\lambda_{m}$ is controlled by adjusting $\alpha$.
 2 Related Works
-Model Merging
-Model merging is a technique that integrates the parameters of multiple models to create a unified model with enhanced or diverse capabilities (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024; Lin et al., 2024).
-Task arithmetic (Ilharco et al., 2022) leverages task vectors for model merging through arithmetic operations, incorporating a predefined scaling term to weight the contribution of different models.
-Fisher Merging (Matena and Raffel, 2022) performs parameter fusion by applying weights derived from the Fisher information matrix (Fisher, 1922), resulting in more precise parameter integration.
-TIES-Merging (Yadav et al., 2023) addresses task conflicts by removing low-magnitude parameters, resolving sign disagreements, and merging only the parameters that align with the final agreed-upon sign.
-In (Yu et al., 2024), it is found that LLMs can enhance their capabilities through model merging.
-Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al., 2022), significantly improving the performance of various model merging techniques.
-DELLA (Deep et al., 2024) is a novel model merging technique that integrates a new pruning strategy called MAGPRUNE, which samples delta parameters based on their magnitudes.
+2.1 Model Merging
+Model merging is a technique that integrates the parameters of multiple models to create a unified model with enhanced or diverse capabilities (Wortsman et al.
+2022; Ilharco et al.
+2022; Matena and Raffel 2022; Jin et al.
+2022; Yadav et al.
+2023; Yu et al.
+2024; Lin et al.
+2024).
+Task arithmetic (Ilharco et al.
+2022) leverages task vectors for model merging through arithmetic operations, incorporating a predefined scaling term to weight the contribution of different models.
+Fisher Merging (Matena and Raffel 2022) performs parameter fusion by applying weights derived from the Fisher information matrix (Fisher 1922), resulting in more precise parameter integration.
+TIES-Merging (Yadav et al.
+2023) addresses task conflicts by removing low-magnitude parameters, resolving sign disagreements, and merging only the parameters that align with the final agreed-upon sign.
+In (Yu et al.
+2024), it is found that LLMs can enhance their capabilities through model merging.
+Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al.
+2022), significantly improving the performance of various model merging techniques.
+DELLA (Deep, Bhardwaj, and Poria 2024) is a novel model merging technique that integrates a new pruning strategy called MAGPRUNE, which samples delta parameters based on their magnitudes.
 MAGPRUNE demonstrates improvements over existing methods such as DARE and TIES.
@@
 3.1 Mixup Model Merge
-Inspired by the random interpolation strategy of Mixup (Zhang, 2017), we extend the idea from input space to parameter space, proposing Mixup Model Merge (M3).
+Inspired by the random interpolation strategy of Mixup (Zhang 2017), we extend the idea from input space to parameter space, proposing Mixup Model Merge (M3).
 We consider two task-specific fine-tuned language models, denoted as $\theta_{t_{1}}^{\text{SFT}}$ and $\theta_{t_{2}}^{\text{SFT}}$, obtained by independently fine-tuning a shared pretrained base model $\theta_{\text{PRE}}$ on two different tasks $t_{1}$ and $t_{2}$, respectively.
 To explore the continuum of models between these two experts, we define a linear interpolation over their parameters:
-$\theta_{M}=\lambda_{m}\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\theta_{\text{SFT}}^{t_{2}},\quad\lambda_{m}\in(0,1)$
+$\theta_{M}=\lambda_{m}\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\theta_{\text{SFT}}^{t_{2}},\quad\lambda_{m}\in(0,1),$ (1)
 This defines a one-dimensional family of interpolated models.
 Alternatively, the interpolation can be expressed in the delta space, where we defined as the parameter offsets between the task-specific fine-tuned model and the pretrained model:
-$\delta_{t_{i}}=\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}},\quad i\in\{1,2\}$
+$\delta_{t_{i}}=\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}},\quad i\in\{1,2\},$ (2)
 Then the interpolated model can be reformulated as:
-$\theta_{M}=\theta_{\text{PRE}}+\lambda_{m}\delta_{t_{1}}+(1-\lambda_{m})\delta_{t_{2}}$
+$\theta_{M}=\theta_{\text{PRE}}+\lambda_{m}\delta_{t_{1}}+(1-\lambda_{m})\delta_{t_{2}},$ (3)
 In both formulations, the resulting model $\theta_{M}$ lies on the line segment connecting the two task-specific models in parameter space.
 This construction naturally defines a one-dimensional model family:
-$\mathcal{F}_{M}=\{\theta_{M}^{\lambda_{m}}\mid\lambda_{m}\in(0,1)\}$
+$\mathcal{F}_{M}=\{\theta_{M}^{\lambda_{m}}\mid\lambda_{m}\in(0,1)\},$ (4)
 Such a model family allows us to study how task behaviors interact and mix at the parameter level, and serves as the foundation for our proposed Mixup Model Merge ($M^{3}$).
-Linear Interpolation Coefficient
+Linear Interpolation Coefficient.
 A key component of the M3 method is the stochastic sampling of the interpolation coefficient $\lambda_{m}$ from a Beta distribution, i.e., $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$.
@@
 As $\alpha$ approaches 1, the distribution of $\lambda_{m}$ becomes more dispersed, leading to lower exploration efficiency but higher diversity in the merged models explored.
-Comparison with Grid Search
-As shown in Table 1, we compare the performance of M3 with that of grid search.
-Grid search requires manually specifying multiple fixed and discrete interpolation coefficients (e.g., $\lambda=0.1,0.2,\ldots,0.9$).
-While this approach offers fine-grained control, it lacks the ability to explore continuous values.
-Achieving finer resolution (e.g., identifying values beyond the first decimal place) requires a smaller step size, which significantly reduces search efficiency.
-In contrast, our method enables more efficient and diverse exploration of continuous interpolation coefficients.
-For example, our approach can sample values such as $\lambda_{m}=0.62156685$, which are difficult to obtain with coarse grid search.
-Empirically, we find that the model’s performance is often sensitive to the second decimal place of $\lambda$.
-To achieve such granularity, grid search would require a step size of 0.01, resulting in a prohibitively large number of evaluations.
-Our method, by comparison, typically requires only a single pass over a small set of hyperparameter values for $\alpha$ (e.g., $\alpha\in\{0.2,0.4,0.5,1,2,3,5\}$).
-In cases where a suitable $\lambda_{m}$ is not found in the first pass, we perform a few additional samplings under a selected $\alpha$ value—no more than three times in our experiments.
-In Table 2, we report the specific $\lambda_{m}$ values that lead to performance improvements, along with the corresponding distribution shape parameter $\alpha$ and the total number of sampling iterations.
-Method Distribution of $\lambda_{m}$ Continuity of $\lambda_{m}$ Efficiency Diversity Controllability
-Grid Search Fixed discrete values (e.g., 0.1–0.9) No Low Medium High
-M3+Beta(1,1) Uniform over (0,1) Yes Medium High Medium
-M3+Beta(5,5) Concentrated near the center (0.5) Yes High Low High
-M3+Beta(0.2,0.2) Concentrated near the endpoints (0 or 1) Yes High Low High
-Table 1: Comparison of interpolation coefficient $\lambda_{m}$ sampling strategies.
-M3 Implementation
+In Table 1, we report the specific $\lambda_{m}$ values that lead to performance improvements, along with the corresponding distribution shape parameter $\alpha$.
+M3 Implementation.
 M3 is a plug-and-play approach that can be integrated into existing representative model merging methods to enhance the performance of the resulting merged models.
-As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022).
+As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al.
+2022) and Task Arithmetic (Ilharco et al.
+2022).
 The official computation process for Average Merging is described as follows:
-$\theta_{M}=\frac{1}{2}\left(\theta_{\text{SFT}}^{t_{1}}+\theta_{\text{SFT}}^{t_{2}}\right)$ (1)
+$\theta_{M}=\frac{1}{2}\left(\theta_{\text{SFT}}^{t_{1}}+\theta_{\text{SFT}}^{t_{2}}\right),$ (5)
 The official computation process for Task Arithmetic is:
-$\theta_{M}$ $=\theta_{\text{PRE}}+\lambda\cdot(\delta^{t_{1}}+\delta^{t_{2}})$
-$=\theta_{\text{PRE}}+\lambda\cdot\sum_{i=1}^{2}(\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}}),$ (2)
+$\theta_{M}$ $=\theta_{\text{PRE}}+\lambda\cdot(\delta^{t_{1}}+\delta^{t_{2}}),$ (6)
+$=\theta_{\text{PRE}}+\lambda\cdot\sum_{i=1}^{2}(\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}}),$
 $\lambda$ is a scaling term that takes values from the set $\{0.0,0.1,\ldots,1.0\}$.
@@
 where
-$\lambda_{m}\in[0,1],\quad\lambda_{m}\sim\mathrm{Beta}(\alpha,\alpha),$
+$\lambda_{m}\in[0,1],\quad\lambda_{m}\sim\mathrm{Beta}(\alpha,\alpha),$ (9)
 $\lambda_{m}$ controls the interpolation ratio, i.e., the contribution ratio, between the two fine-tuned models. $\alpha$ determines the shape of the Beta distribution.
 Figure 3: The Beta distribution visualization for different $\alpha$ values.
+Merging
+Methods
+Models
+Use M3
+Sampling
+Details
+Instruction
+Following
+Mathematical
+Reasoning
+Code Generating
+Avg.
+$\lambda_{m}$ $\alpha$ AlpacaEval GSM8K MATH HumanEval MBPP
+- LM - - - 45.14 2.20 0.04 36.59 34.00 -
+Math - - - - 64.22 14.02 - - -
+Code - - - - - - 23.78 27.60 -
+Average
+Merging
+LM
+& Math
+No - - 45.28 66.34 13.40 - - 41.67
+Yes 0.35 2 44.40 66.26 13.80 - - 41.48
+LM
+& Code
+No - - 36.60 - - 29.88 32.00 32.82
+[…]
+Math
+& Code
+No - - - 56.17 10.28 8.53 8.20 20.80
+[…]
+Task
+Arithmetic
+LM
+& Math
+No - - 45.78 66.34 13.40 - - 41.84
+Yes 0.34 2 41.65 66.34 13.74 - - 40.58
+LM
+& Code
+No - - 44.64 - - 32.93 33.60 37.06
+Yes 0.99 0.4 46.64 - - 35.37 33.80 38.60
+Math
+& Code
+No - - - 64.67 13.98 8.54 8.60 23.95
+[…]
+TIES-
+Merging
+LM
+& Math
+No - - 38.63 14.56 2.12 - - 18.44
+[…]
+LM
+& Code
+No - - 41.85 - - 0.0 0.0 13.95
+[…]
+Math
+& Code
+No - - - 64.67 13.68 9.15 22.60 27.57
+Yes 0.63 3 - 64.75 14.16 9.76 21.4 27.52
+DARE
+LM
+& Math
+No - - 49.00 66.64 13.2 - - 42.95
+Yes 0.38 0.4 44.90 67.32 13.74 - - 41.98
+LM
+& Code
+No - - 41.47 - - 35.98 33.00 36.82
+[…]
+Math
+& Code
+No - - - 65.05 10.37 9.15 9.80 23.59
+[…]
+Table 1: Comparison of Merged Model Performance Before and After Applying M3.
+We use three task-specific LLMs: WizardLM-13B (LM), WizardMath-13B (Math), and llama-2-13b-codealpaca (Code).
+Bold indicates the best results for each dataset, underline highlights improvements with M3, and $\uparrow$ denotes significant gains after introducing M3.
 3.2 Theoretical Analysis
-According to the theoretical and empirical findings of Neyshabur et al.
-(2020), fine-tuning models from the same pretrained initialization can guide them into a shared, flat, and connected low-loss basin.
+According to the theoretical and empirical findings of Neyshabur, Sedghi, and Zhang (2020), fine-tuning models from the same pretrained initialization can guide them into a shared, flat, and connected low-loss basin.
 As a result, the linear interpolation path between such fine-tuned models remains within this low-loss basin and does not encounter significant performance barriers.
@@
 Define a function representing the change of the merged model’s $i$-th parameter relative to the pretrained model:
-$f_{i}(\lambda_{m})=\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i}.$ (6)
+$f_{i}(\lambda_{m})=\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i},$ (11)
 If $\delta_{t1,i}>0$ and $\delta_{t2,i}<0$, the delta parameters are in conflict and may partially cancel each other. In particular, when
@@
 4.1 Experimental Setup
-Task-Specific Fine-Tuned LLMs and Datasets
+Task-Specific Fine-Tuned LLMs and Datasets.
 Following the experimental setup given in Yu et al.
-(2024), we select three task-specific fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), all of which use Llama-2-13b (Touvron et al., 2023) as the pre-trained backbone.
+(2024), we select three task-specific fine-tuned LLMs: WizardLM-13B (Xu et al.
+2024), WizardMath-13B (Luo et al.
+2023), and llama-2-13b-code-alpaca (Chaudhary 2023), all of which use Llama-2-13b (Touvron et al.
+2023) as the pre-trained backbone.
 These models are respectively designed for instruction-following, mathematical reasoning, and code generation tasks.
-To evaluate the instruction-following task we use AlpacaEval (Li et al., 2023).
-For testing mathematical reasoning task, we employ GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
-For estimating the performance of code-generating task, we use HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
-More details of these LLMs and datasets can be found in Appendix A.1.
-Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness
-To evaluate OOD robustness, we evaluate merged models including math & code, LM & math, and LM & code on instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al., 2024), respectively.
-More details on OOD benchmarks are given in Appendix A.3.
-We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al., 2024) to evaluate the robustness of LLMs against adversarial prompts.
-Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al., 2018), BERTAttack (word-level) (Li et al., 2020), and StressTest (sentence-level) (Naik et al., 2018).
-The evaluation is conducted on two datasets supported by PromptBench: SST2 (sentiment analysis) (Socher et al., 2013) and CoLA (grammatical correctness) (Warstadt, 2019).
-For more details on PromptBench and attack methods, please refer to Appendix A.4.
-Evaluation Metrics
+To evaluate the instruction-following task we use AlpacaEval (Li et al.
+2023).
+For testing mathematical reasoning task, we employ GSM8K (Cobbe et al.
+2021) and MATH (Hendrycks et al.
+2021).
+For estimating the performance of code-generating task, we use HumanEval (Chen et al.
+2021) and MBPP (Austin et al.
+2021).
+More details of these LLMs and datasets can be found in Appendix A.
+Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness.
+To evaluate OOD robustness, we evaluate merged models including math & code, LM & math, and LM & code on instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al.
+2024), respectively.
+More details on OOD benchmarks are given in Appendix B.
+We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al.
+2024) to evaluate the robustness of LLMs against adversarial prompts.
+Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al.
+2018), BERTAttack (word-level) (Li et al.
+2020), and StressTest (sentence-level) (Naik et al.
+2018).
+The evaluation is conducted on two datasets supported by PromptBench: SST2 (sentiment analysis) (Socher et al.
+2013) and CoLA (grammatical correctness) (Warstadt 2019).
+For more details on PromptBench and attack methods, please refer to Appendix C.
+Evaluation Metrics.
 We calculate win rate for AlpacaEval and LiveBench-Instruction, zero-shot accuracy for GSM8K and MATH, pass@1 for HumanEval, MBPP and LiveBench-Coding, Matthews correlation coefficient (MCC) for CoLA, accuracy for SST2, and zero-shot accuracy for LiveBench-TypoFixing.
-Implementation Details
+Implementation Details.
 In each distinct merging experiment—where “distinct” refers to variations in either the task-specific LLMs or the model merging methods—we perform a full sweep over the sole hyperparameter of M3, $\alpha$, using the set $\{0.2,0.4,0.5,1,2,3,5\}$.
@@
 (2024).
-For a detailed description of the hyperparameter settings used in the representative existing model merging methods, please refer to Appendix A.2.
+For a detailed description of the hyperparameter settings used in the representative existing model merging methods, please refer to Appendix D.
 All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
+Model Dataset
+Use
+Mixup
+Use
+Attack
+Metric (%) PDR (%)
+Math
+Code
+SST2 No No 57.68 38.97
+Yes 35.21
+Yes No 86.24 35.77
+Yes 55.39
+CoLA No No 45.54 98.53
+Yes 0.67
+Yes No 71.72 6.42
+Yes 67.11
+LM
+Math
+SST2 No No 92.78 29.05
+Yes 65.83
+Yes No 91.28 34.55
+Yes 59.75
+CoLA No No 79.19 8.84
+Yes 72.20
+Yes No 80.54 4.52
+Yes 76.89
+LM
+Code
+SST2 No No 10.55 38.04
+Yes 6.54
+Yes No 73.17 7.68
+Yes 67.55
+CoLA No No 74.21 47.42
+Yes 39.02
+Yes No 74.78 31.67
+Yes 51.10
+Table 2: Adversarial robustness of merged models on SST2 and CoLA with StressTest prompt attacks. Best and second-best results are marked in bold and underlined, respectively.
+Merging
+Methods
+Models
+Use M3
+Use
+DARE
+Instruction
+Following
+Mathematical
+Reasoning
+Code Generating
+AlpacaEval GSM8K MATH HumanEval MBPP
+Average
+Merging
+LM
+& Math
+Yes No 44.40 66.26 13.80 - -
+No Yes 44.22 66.57 12.96 - -
+Yes Yes 43.53 66.57 14.12 - -
+LM
+& Code
+Yes No 43.91 - - 37.20 34.40
+No Yes 38.81 - - 31.71 32.40
+Yes Yes 40.31 - - 36.59 37.00
+Math
+& Code
+[…]
+No Yes - 56.18 10.28 6.10 7.80
+Yes Yes - 64.97 13.54 9.76 21.20
+Task
+Arithmetic
+LM
+& Math
+Yes No 41.65 66.34 13.74 - -
+No Yes 49.00 66.64 13.02 - -
+Yes Yes 44.90 67.32 13.74 - -
+LM
+& Code
+Yes No 46.64 - - 35.37 33.80
+No Yes 41.47 - - 35.98 33.00
+Yes Yes 45.20 - - 35.98 35.20
+Math
+& Code
+Yes No - 63.53 13.94 7.93 19.00
+No Yes - 65.05 13.96 10.37 9.80
+Yes Yes - 65.13 14.32 8.54 18.00
+TIES-
+Merging
+LM
+& Math
+Yes No 38.73 18.57 2.48 - -
+No Yes 37.92 18.04 2.34 - -
+Yes Yes 39.93 19.26 2.82 - -
+LM
+& Code
+Yes No 44.96 - - 25.61 30.80
+No Yes 43.13 - - 0.0 0.0
+Yes Yes 45.65 - - 26.83 33.20
+Math
+& Code
+Yes No - 64.75 14.16 9.76 21.4
+No Yes - 64.82 13.88 10.37 23.60
+Yes Yes - 64.75 14.78 9.15 19.60
+Table 3: Comparison of our method M3 and DARE. The best and second-best results are marked in bold and underlined fonts.
 4.2 Merging Task-Specific Fine-Tuned LLMs
-Merging Methods Models Use M3 (Ours) Sampling Details Instruction Following Mathematical Reasoning Code Generating Avg.
-$\lambda_{m}$ $\alpha$ AlpacaEval GSM8K MATH HumanEval MBPP
-- LM - - - 45.14 2.20 0.04 36.59 34.00 -
-Math - - - - 64.22 14.02 - - -
-Code - - - - - - 23.78 27.60 -
-Average Merging LM & Math No - - 45.28 66.34 13.40 - - 41.67
-Yes 0.35 2 44.40 66.26 13.80 - - 41.48
-LM & Code No - - 36.60 - - 29.88 32.00 32.82
-Yes 0.84 2 43.91$\uparrow$ - - 37.20$\uparrow$ 34.40 38.50$\uparrow$
-Math & Code No - - - 56.17 10.28 8.53 8.20 20.80
-Yes 0.92 0.4 - 63.61$\uparrow$ 14.02$\uparrow$ 8.54 19.20$\uparrow$ 26.34$\uparrow$
-Task Arithmetic LM & Math No - - 45.78 66.34 13.40 - - 41.84
-Yes 0.34 2 41.65 66.34 13.74 - - 40.58
-LM & Code No - - 44.64 - - 32.93 33.60 37.06
-Yes 0.99 0.4 46.64 - - 35.37 33.80 38.60
-Math & Code No - - - 64.67 13.98 8.54 8.60 23.95
-Yes 0.98 0.5 - 63.53 13.94 7.93 19.00$\uparrow$ 26.1
-TIES- Merging LM & Math No - - 38.63 14.56 2.12 - - 18.44
-Yes 0.62 1 38.73 18.57$\uparrow$ 2.48 - - 19.92
-LM & Code No - - 41.85 - - 0.0 0.0 13.95
-Yes 0.84 0.5 44.96 - - 25.61$\uparrow$ 30.80$\uparrow$ 33.79$\uparrow$
-Math & Code No - - - 64.67 13.68 9.15 22.60 27.57
-Yes 0.63 3 - 64.75 14.16 9.76 21.4 27.52
-DARE LM & Math No - - 49.00 66.64 13.2 - - 42.95
-Yes 0.38 0.4 44.90 67.32 13.74 - - 41.98
-LM & Code No - - 41.47 - - 35.98 33.00 36.82
-Yes 0.99 0.5 45.20$\uparrow$ - - 35.98 35.20 38.79
-Math & Code No - - - 65.05 10.37 9.15 9.80 23.59
-Yes 0.98 0.5 - 65.13 14.32$\uparrow$ 8.54 18.00$\uparrow$ 26.50$\uparrow$
-Table 2: Comparison of Merged Model Performance Before and After Applying M3.
-We use three task-specific LLMs: WizardLM-13B (LM), WizardMath-13B (Math), and llama-2-13b-codealpaca (Code).
-Bold indicates the best results for each dataset, underline highlights improvements with M3, and $\uparrow$ denotes significant gains after introducing M3.
 Our proposed method, M3, is a plug-and-play approach.
@@
 The results is presented in Table 1.
-Performance Gains of M3 Across Merging Methods and Tasks
+Performance Gains of M3 Across Merging Methods and Tasks.
 As shown in Table 1, M3 generally enhances existing model merging methods, achieving over 10-point improvements on multiple datasets and up to 30 points on one.
@@
 This suggests that M3 enables the discovery of interpolated models along linear paths between two models that outperform both endpoints, providing support for the theoretical insights discussed in Section 3.2.
-In addition, we also compare our proposed method M3 with existing model merging methods. The details are in Appendix G.
-Mitigating Suboptimal Fine-Tuning with M3
+In addition, we also compare our proposed method M3 with existing model merging methods. The details are in Appendix E.
+Mitigating Suboptimal Fine-Tuning with M3.
 Yu et al.
@@
 In other words, M3 significantly mitigates the negative impact of insufficient fine-tuning on merging quality, further highlighting the critical importance of task-specific LLMs’ contribution ratio in determining the performance of the merged model.
+Figure 4: Performance of merged models (Math & Code, LM & Math, and LM & Code) using three model merging methods (Average Merging, Task Arithmetic, and TIES-Merging) on OOD datasets.
 4.3 Model Robustness
-Out-of-distribution robustness
-Figure 4: Performance of merged models (Math & Code, LM & Math, and LM & Code) using three model merging methods (Average Merging, Task Arithmetic, and TIES-Merging) on OOD datasets.
-To ensure that the evaluation datasets are as representative as possible of OOD data, we select datasets with sufficiently recent release dates and ensure they cover domains that fine-tuned LLMs have not been specifically trained on.
-Consequently, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing.
-As illustrated in Figure 4, M3 consistently enhances the performance of merged models—Math & Code, LM & Math, and LM & Code—on OOD datasets.
-Specifically, when Task Arithmetic is combined with M3, the Math & Code model demonstrates a 1.9% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 1.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model shows a significant 6% boost in accuracy on LiveBench-TypoFixing.
-Similarly, when Average Merging is combined with M3, the Math & Code model attains a 1.5% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 0.7% increase in pass@1 on LiveBench-Coding, and the LM & Code model exhibits a 4% enhancement in accuracy on LiveBench-TypoFixing.
-Finally, when TIES-Merging is applied alongside M3, the Math & Code model achieves a 1.1% improvement in win rate on LiveBench-Instruction, the LM & Math model records a 0.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model demonstrates a remarkable 14% improvement in accuracy on LiveBench-TypoFixing.
-Adversarial robustness
-Model Dataset Use Mixup Use Attack Metric (%) PDR (%)
-Math & Code SST2 No No 57.68 38.97
-Yes 35.21
-Yes No 86.24 35.77
-Yes 55.39
-CoLA No No 45.54 98.53
-Yes 0.67
-Yes No 71.72 6.42
-Yes 67.11
-LM & Math SST2 No No 92.78 29.05
-Yes 65.83
-Yes No 91.28 34.55
-Yes 59.75
-CoLA No No 79.19 8.84
-Yes 72.20
-Yes No 80.54 4.52
-Yes 76.89
-LM & Code SST2 No No 10.55 38.04
-Yes 6.54
-Yes No 73.17 7.68
-Yes 67.55
-CoLA No No 74.21 47.42
-Yes 39.02
-Yes No 74.78 31.67
-Yes 51.10
-Table 3: Adversarial Robustness of Merged Models on the SST2 and CoLA Datasets when Executing StressTest prompt attack method. The best and second-best results are marked in bold and underlined fonts.
-We employ three Prompt Attack Methods supported by the promptbench codebase (DeepWordBug, BERTAttack, and StressTest) (Zhu et al., 2024) to evaluate the adversarial robustness of three merged models (Math & Code, LM & Math, and LM & Code) obtained through the task arithmetic method.
-To balance experimental effectiveness with computational efficiency, we randomly selected the positions of the three attacked words in the prompts when executing the DeepWordBug and BERTAttack attacks.
-Adversarial robustness is assessed using the Performance Drop Rate (PDR) (Zhu et al., 2023), where a lower PDR indicates stronger robustness.
-Further details on PDR can be found in Appendix D.
-As shown in Table 3, M3 improves the adversarial robustness of the merged models in most cases with the StressTest Prompt Attack Method.
-For example, with M3, the PDR of Math & Code decreased by 3.2% on the SST2 dataset and by 92.12% on the CoLA dataset, while the PDR of LM & Code decreased by 30.36% on SST2 and by 15.75% on CoLA.
-Furthermore, in most cases, M3 not only improves the adversarial robustness of the merged models but also enhances their performance metrics (accuracy and MCC) on the SST2 and CoLA datasets.
-Specifically, with M3, Math & Code demonstrates a 28.56% increase in accuracy on SST2 and a 26.18% increase in MCC on CoLA, while LM & Code achieves a 62.62% increase in accuracy on SST2.
-These results show that M3 effectively enhances both the adversarial robustness and the performance of the merged models in sentiment analysis and grammar correctness tasks.
-Detailed experimental results for the remaining Prompt Attack Methods (DeepWordBug and BERTAttack) are presented in Appendix B.
+Out-of-Distribution Robustness.
+To ensure the evaluation datasets reflect true OOD scenarios, we select recently released datasets from domains not seen during fine-tuning.
+Accordingly, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing.
+As shown in Figure 4, M3 consistently improves OOD performance across all model pairs and merging methods.
+For example, under Task Arithmetic, M3 yields gains of 1.9%, 1.6%, and 6% on the three respective tasks.
+Similar improvements are observed with Average Merging (1.5%, 0.7%, 4%) and TIES-Merging (1.1%, 0.6%, 14%).
+These results highlight M3’s effectiveness in enhancing OOD generalization.
+Adversarial robustness.
+We evaluate the adversarial robustness of three merged models (Math & Code, LM & Math, and LM & Code) using three prompt attack methods from PromptBench (Zhu et al.
+2024): DeepWordBug, BERTAttack, and StressTest.
+For DeepWordBug and BERTAttack, we randomly select three word positions per prompt.
+Robustness is measured by Performance Drop Rate (PDR) (Zhu et al.
+2023), with lower PDR indicating stronger robustness (see Appendix F for details).
+As shown in Table 2, M3 notably improves robustness under StressTest attacks.
+For instance, it reduces PDR by 3.2% (SST2) and 92.12% (CoLA) for Math & Code, and by 30.36% (SST2) and 15.75% (CoLA) for LM & Code.
+Additionally, M3 improves performance metrics such as accuracy and MCC—e.g., accuracy for LM & Code on SST2 increases by 62.62%.
+These results indicate that M3 enhances both robustness and task performance.
+Results for DeepWordBug and BERTAttack are provided in Appendix G.
 4.4 Comparison of Mixup Model Merge and DARE
-DARE is a model sparsification method proposed by (Yu et al., 2024), with a more detailed introduction provided in Appendix E.
-This section takes DARE as an example to discuss the relationship between the existing model sparsification method and M3.
-We integrated M3 and DARE with three existing model merging techniques—Average Merging, Task Arithmetic, and TIES-Merging—to compare these two plug-and-play methods.
-Detailed experimental results are provided in Appendix F.
-Additionally, in the DARE method, the drop rate hyperparameter is set to 0.2.
-From Table 8, it can be concluded that M3 generally outperforms DARE, exhibiting significantly superior performance on certain datasets.
-For instance, the Math & Code model achieves a pass@1 score of 9.8% on the MBPP dataset when combined with DARE, while this score increases to 19% when combined with M3.
-As shown in Table 8, Combining DARE and M3 generally results in better model merging performance.
-This suggests that M3 and DARE can complement each other.
-Moreover, in some cases, M3 alone can deliver the best results, while DARE alone only achieves the best performance in very few cases, further demonstrating the superiority of M3.
+DARE is a model sparsification method proposed by Yu et al.
+(2024) (see Appendix H for details).
+We compare M3 and DARE by integrating both into three representative model merging methods: Average Merging, Task Arithmetic, and TIES-Merging.
+The drop rate for DARE is fixed at 0.2.
+As shown in Table 3, M3 generally outperforms DARE and achieves significantly better results on several datasets.
+For example, on MBPP, the pass@1 score of the Math & Code model increases from 9.8% with DARE to 19% with M3.
+Moreover, combining M3 with DARE typically leads to even stronger performance, suggesting they are complementary.
+Notably, M3 alone often yields the best results, while DARE alone rarely does, further highlighting the effectiveness of M3.
 5 Conclusion
@@
 By tuning the Beta distribution parameters, M3 balances efficiency and diversity in exploring contribution ratios.
-6 Limitations
-Despite its advantages, the M3 method has several limitations: (1) Our method is currently limited to pairwise model merging, as random linear interpolation is typically defined between two entities.
-Extending it to merge three or more models requires new strategies for sampling multi-way interpolation coefficients, which introduces additional complexity and remains an open direction for future work.
-(2) This study focuses on merging homogeneous models fine-tuned from the same pretrained backbone.
-Applying M3 to heterogeneous models with different architectures or fine-tuning paths poses new challenges and will be explored in future work.
-(3) Our method may also be extended to broader applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) models to reduce the alignment tax (FINE-TUNING, ).
-Appendix A Detailed Experimental Settings
-A.1 Task-Specific Fine-Tuned LLMs and Datasets Details
+This simple and efficient method offers new insights for optimizing model merging strategies.
+In the future, M3 may also be extended to broader applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) models (Ouyang et al.
+2022), potentially reducing the alignment tax associated with post-hoc alignment techniques like fine-tuning.
+Appendix A Task-Specific Fine-Tuned LLMs and Datasets Details
 We conduct model merging experiments using three task-specific LLMs fine-tuned from Llama-2-13b:
   • WizardLM-13B is an instruction-following model based on Llama-2-13b, designed to improve open-domain instruction-following.
-Using the Evol-Instruct method (Xu et al., 2024), it generates high-complexity instruction data to reduce human annotation and enhance generalization.
+Using the Evol-Instruct method (Xu et al.
+2024), it generates high-complexity instruction data to reduce human annotation and enhance generalization.
 The model undergoes supervised fine-tuning with AI-generated data, followed by refinement via RLHF.
 Evaluation results show that Evol-Instruct-generated instructions outperform human-written ones, and WizardLM-13B surpasses ChatGPT in high-complexity tasks.
-In GPT-4 automated evaluation, it achieves over 90% of ChatGPT’s performance in 17 out of 29 tasks, demonstrating the effectiveness of AI-evolved instruction fine-tuning (Xu et al., 2024).
-  • WizardMath-13B, optimized from Llama-2-13b, is designed for mathematical reasoning and enhances Chain-of-Thought (CoT) (Wei et al., 2022) capabilities.
+In GPT-4 automated evaluation, it achieves over 90% of ChatGPT’s performance in 17 out of 29 tasks, demonstrating the effectiveness of AI-evolved instruction fine-tuning (Xu et al.
+2024).
+  • WizardMath-13B, optimized from Llama-2-13b, is designed for mathematical reasoning and enhances Chain-of-Thought (CoT) (Wei et al.
+2022) capabilities.
 It uses Reinforcement Learning from Evol-Instruct Feedback to evolve math tasks and improve reasoning.
@@
   • llama-2-13b-code-alpaca is a code generation model fine-tuned from Llama-2-13b, designed to enhance code understanding and generation.
-It follows the same training approach as Stanford Alpaca (Taori et al., 2023) but focuses on code-related tasks.
-The model is fine-tuned with 20K instruction-following code samples generated using the Self-Instruct method (Wang et al., 2022).
+It follows the same training approach as Stanford Alpaca (Taori et al.
+2023) but focuses on code-related tasks.
+The model is fine-tuned with 20K instruction-following code samples generated using the Self-Instruct method (Wang et al.
+2022).
 However, as it has not undergone safety fine-tuning, caution is required when using it in production environments.
 We use one dataset to evaluate the instruction-following task:
-  • AlpacaEval (Li et al., 2023) is an LLM-based automated evaluation metric that assesses model performance by testing on a fixed set of 805 instructions and computing the win rate of the evaluated model against a baseline.
+  • AlpacaEval (Li et al.
+2023) is an LLM-based automated evaluation metric that assesses model performance by testing on a fixed set of 805 instructions and computing the win rate of the evaluated model against a baseline.
 The evaluation process involves an LLM-based evaluator that compares the responses and determines the probability of preferring the evaluated model.
-In this paper, we use AlpacaEval 2.0 (Dubois et al., 2024).
+In this paper, we use AlpacaEval 2.0 (Dubois et al.
+2024).
 To reduce costs, we use chatgpt_fn for evaluation.
@@
 It consists of 7.5K training problems and 1K test problems.
-In this paper, we use the 1K test set for evaluation (Cobbe et al., 2021).
+In this paper, we use the 1K test set for evaluation (Cobbe et al.
+2021).
   • MATH is a dataset containing 12,500 competition-level math problems, designed to evaluate and enhance the problem-solving abilities of machine learning models.
 It consists of 7,500 training problems and 5,000 test problems.
-We use the 5,000 test set for evaluation (Hendrycks et al., 2021).
+We use the 5,000 test set for evaluation (Hendrycks et al.
+2021).
 We used two dataset to evaluate the code generation task:
@@
 Each problem includes a function signature, docstring, function body, and unit tests.
-The dataset tests models’ language comprehension, reasoning, and algorithmic abilities (Chen et al., 2021).
+The dataset tests models’ language comprehension, reasoning, and algorithmic abilities (Chen et al.
+2021).
   • MBPP is a dataset containing 974 programming problems designed to evaluate a model’s ability to synthesize Python programs from natural language descriptions.
 The problems range from basic numerical operations to more complex tasks involving list and string processing.
-The test set consists of 500 problems, which are used for evaluation in this paper (Austin et al., 2021).
-A.2 Hyperparameter Setting Details in Model Merging Methods
+The test set consists of 500 problems, which are used for evaluation in this paper (Austin et al.
+2021).
+Appendix B Out-of-Distribution Dataset Selection Details
+LiveBench (White et al.
+2024) is a dynamic benchmark for large language models, featuring frequently updated questions and diverse tasks.
+To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench, respectively, deliberately avoiding the fine-tuning domains of the merged fine-tuned models.
+These tasks were released after November 2023, whereas WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca were all introduced earlier.
+Furthermore, their shared Llama-2-13b backbone was trained on data only up to July 2023.
+Consequently, these factors collectively ensure that the evaluation remains OOD in the temporal dimension.
+When assessing the OOD robustness of LM & Code using the Language Comprehension category in LiveBench, only the typo-fixing task is considered.
+This decision is based on the fact that LiveBench is highly challenging, and the merged model performs poorly on other tasks in this category, with accuracy close to zero, rendering the evaluation results inconclusive and uninformative.
+Finally, we acknowledge the limitations of these datasets.
+For large models like Llama-2-13b, identifying truly OOD datasets is difficult, as their training data likely covers similar distributions.
+These datasets are better described as ”out-of-example”, representing instances not explicitly seen during training.
+As discussed in (Wang et al.
+2023), distribution shifts can occur across domains and time.
+While Llama-2-13b may have been trained on datasets for tasks like instruction-following, coding, and language comprehension, the datasets we selected remain valuable for OOD evaluation by capturing temporal shifts, providing insights into robustness over time.
+Appendix C Adversarial Robustness Evaluation Experiments Setting Details
+PromptBench (Zhu et al.
+2024) is a unified library designed for evaluating LLMs, providing a standardized and extensible framework.
+It includes several key components such as prompt construction, prompt engineering, dataset and model loading, adversarial prompt attacks, dynamic evaluation protocols, and analysis tools.
+We use the Adversarial Prompt Attacks module in PromptBench aims to evaluate the robustness of LLMs against adversarial prompts.
+We employ three methods to perform adversarial attacks on prompts to evaluate the adversarial robustness of the merged models: DeepWordBug (Gao et al.
+2018), BERTAttack (Li et al.
+2020), and StressTest (Naik et al.
+2018), representing Character-level, Word-level, and Sentence-level attacks, respectively.
+  • DeepWordBug introduces subtle character-level perturbations (e.g., adding, deleting, or replacing characters) to words in text to deceive language models.
+It aims to evaluate a model’s robustness against small typographical errors that may alter the model’s performance without being easily detected.
+  • BERTAttack manipulates text at the word level by replacing words with contextually similar synonyms to mislead large language models.
+This method tests the model’s ability to maintain accuracy despite small lexical changes that might alter the meaning of the input.
+  • StressTest appends irrelevant or redundant sentences to the end of a prompt to distract and confuse language models.
+It assesses the model’s ability to handle extraneous information and maintain accuracy when faced with unnecessary distractions.
+(a) The operational steps of TIES-Merging.
+(b) After introducing M3, the Disjoint Merge step in the TIES-Merging procedure.
+Figure 5: The difference between M3 and the original TIES-Merging is that, in the Disjoint Merge step, when two task vectors are retained for a given parameter, the mean of the task vectors is replaced by a random linear interpolation, while the other operations remain unchanged.
+The evaluation is conducted on the Sentiment Analysis dataset (SST2 (Socher et al. 2013)) and the Grammar Correctness dataset (CoLA (Warstadt 2019)):
+  • SST2 (Socher et al. 2013): A sentiment analysis dataset designed to assess whether a given sentence conveys a positive or negative sentiment.
+  • CoLA (Warstadt 2019): A dataset for grammar correctness, where the model must determine whether a sentence is grammatically acceptable.
+Appendix D Hyperparameter Setting Details in Model Merging Methods
 Table 4 presents the hyperparameter search ranges for the model merging methods.
@@
 These settings are further applied to model merging experiments involving M3 and DARE.
-Merging Method Model Hyperparameter Values
-TIES-Merging LM & Math scaling_term=0.5, retain_ratio=0.5
-LM & Code scaling_term=1.0, retain_ratio=0.7
-Math & Code scaling_term=1.0, retain_ratio=0.5
+Merging Method
+Model
+Hyperparameter Values
+TIES-Merging
+LM & Math
+scaling_term=0.5, retain_ratio=0.5
+LM & Code
+scaling_term=1.0, retain_ratio=0.7
+Math & Code
+scaling_term=1.0, retain_ratio=0.5
 Table 5: Hyperparameter settings in TIES-Merging.
-A.3 Out-of-Distribution Dataset Selection Details
-LiveBench (White et al., 2024) is a dynamic benchmark for large language models, featuring frequently updated questions and diverse tasks.
-To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench, respectively, deliberately avoiding the fine-tuning domains of the merged fine-tuned models.
-These tasks were released after November 2023, whereas WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca were all introduced earlier.
-Furthermore, their shared Llama-2-13b backbone was trained on data only up to July 2023.
-Consequently, these factors collectively ensure that the evaluation remains OOD in the temporal dimension.
-When assessing the OOD robustness of LM & Code using the Language Comprehension category in LiveBench, only the typo-fixing task is considered.
-This decision is based on the fact that LiveBench is highly challenging, and the merged model performs poorly on other tasks in this category, with accuracy close to zero, rendering the evaluation results inconclusive and uninformative.
-Finally, we acknowledge the limitations of these datasets.
-For large models like Llama-2-13b, identifying truly OOD datasets is difficult, as their training data likely covers similar distributions.
-These datasets are better described as "out-of-example", representing instances not explicitly seen during training.
-As discussed in (Wang et al., 2023), distribution shifts can occur across domains and time.
-While Llama-2-13b may have been trained on datasets for tasks like instruction-following, coding, and language comprehension, the datasets we selected remain valuable for OOD evaluation by capturing temporal shifts, providing insights into robustness over time.
-A.4 Adversarial Robustness Evaluation Experiments Setting Details
-PromptBench (Zhu et al., 2024) is a unified library designed for evaluating LLMs, providing a standardized and extensible framework.
-It includes several key components such as prompt construction, prompt engineering, dataset and model loading, adversarial prompt attacks, dynamic evaluation protocols, and analysis tools.
-We use the Adversarial Prompt Attacks module in PromptBench aims to evaluate the robustness of LLMs against adversarial prompts.
-We employ three methods to perform adversarial attacks on prompts to evaluate the adversarial robustness of the merged models: DeepWordBug (Gao et al., 2018), BERTAttack (Li et al., 2020), and StressTest (Naik et al., 2018), representing Character-level, Word-level, and Sentence-level attacks, respectively.
-  • DeepWordBug introduces subtle character-level perturbations (e.g., adding, deleting, or replacing characters) to words in text to deceive language models.
-It aims to evaluate a model’s robustness against small typographical errors that may alter the model’s performance without being easily detected.
-  • BERTAttack manipulates text at the word level by replacing words with contextually similar synonyms to mislead large language models.
-This method tests the model’s ability to maintain accuracy despite small lexical changes that might alter the meaning of the input.
-  • StressTest appends irrelevant or redundant sentences to the end of a prompt to distract and confuse language models.
-It assesses the model’s ability to handle extraneous information and maintain accuracy when faced with unnecessary distractions.
-(a) The operational steps of TIES-Merging.
-(b) After introducing M3, the Disjoint Merge step in the TIES-Merging procedure.
-Figure 5: The difference between M3 and the original TIES-Merging is that, in the Disjoint Merge step, when two task vectors are retained for a given parameter, the mean of the task vectors is replaced by a random linear interpolation, while the other operations remain unchanged.
-The evaluation is conducted on the Sentiment Analysis dataset (SST2 (Socher et al., 2013)) and the Grammar Correctness dataset (CoLA (Warstadt, 2019)):
-  • SST2 (Socher et al., 2013): A sentiment analysis dataset designed to assess whether a given sentence conveys a positive or negative sentiment.
-  • CoLA (Warstadt, 2019): A dataset for grammar correctness, where the model must determine whether a sentence is grammatically acceptable.
-Appendix B Additional Experimental Results on Adversarial Robustness
+Appendix E Comparison of M3 and Existing Model Merging Approaches
+Models
+Merging Methods
+Instruction
+Following
+Mathematical
+Reasoning
+Code
+Generating
+Avg.
+AlpacaEval GSM8K MBPP
+LM
+& Math
+Average Merging
+45.28 66.34 - 55.81
+Task Arithmetic
+45.78 66.34 - 56.06
+TIES-Merging
+38.63 14.56 - 26.60
+DARE 49.00 66.64 - 57.82
+DELLA 44.03 48.45 - 46.24
+M3 (Ours) 44.90 67.32 - 56.11
+LM
+& Code
+Average Merging
+36.60 - 32.00 34.3
+Task Arithmetic
+44.64 - 33.60 39.12
+TIES-Merging
+41.85 - 0.0 20.93
+DARE 41.47 - 33.00 37.24
+DELLA 40.99 - 35.00 38.00
+M3 (Ours) 45.20 - 35.20 40.2
+Math
+& Code
+Average Merging
+- 56.17 8.20 32.19
+Task Arithmetic
+- 64.67 8.60 36.64
+TIES-Merging
+- 64.67 22.60 43.64
+DARE - 65.05 9.80 37.43
+DELLA - 63.08 20.2 41.64
+M3 (Ours) - 64.75 21.4 43.10
+Table 6: Comparison of our method M3 and existing model merging approaches. The best and second-best results are marked in bold and underlined fonts.
+For each baseline method—Average Merging, Task Arithmetic, TIES-Merging, and DARE—we combine it with our proposed M3 method and select the best-performing combination for each merged model (LM & Math, LM & Code and Math & Code).
+These optimal results are reported as M3 (Ours) in Table 6.
+As presented in Table 6, our proposed method M3 achieves the best overall performance across the three evaluated tasks.
+Importantly, M3 attains these improvements without involving complex architectural designs or intensive computations; it merely adjusts the contribution ratios of constituent models to realize substantial gains, thereby outperforming the DELLA approach.
+DELLA advances model sparsification techniques by building upon TIES-Merging and DARE, introducing a novel pruning strategy termed MAGPRUNE, which selects delta parameters based on their magnitudes.
+This strategy has inspired us to investigate new directions for enhancing the performance of merged models.
+Appendix F Performance Drop Rate (PDR) for Adversarial Robustness
+The adversarial robustness is evaluated using the Performance Drop Rate (PDR) (Zhu et al. 2023), which is defined as follows:
+$\text{PDR}=\frac{\text{Metric}_{\text{no attack}}-\text{Metric}_{\text{attack}}}{\text{Metric}_{\text{no attack}}},$ (13)
+where $\text{Metric}_{\text{no attack}}$ denotes the performance metric without any prompt attack, and $\text{Metric}_{\text{attack}}$ represents the performance metric under the prompt attack.
+A smaller PDR indicates stronger adversarial defense against prompt attacks, implying better adversarial robustness.
+Appendix G Additional Experimental Results on Adversarial Robustness
 Model Dataset Use Mixup Use Attack Metric (%) PDR (%)
@@
 Table 8 presents the detailed experimental results of the adversarial robustness of merged models on the SST2 and CoLA datasets applying the BERTAttack prompt attack method.
-Appendix C Integrating M3 into the TIES-Merging Model Merging Method
-Figure 5 shows the specific implementation approach to incorporating M3 into TIES-Merging.
-After the steps of trimming parameters with lower magnitudes and resolving sign disagreements, the two models to be merged are denoted as $M_{1}$ and $M_{2}$.
-During the M3 process, only the parameters that are preserved in both $M_{1}$ and $M_{2}$ are interpolated according to the model merging hyperparameter $\lambda_{m}$ to obtain the merged parameters.
-For parameters that are preserved in only one of the models, no interpolation is performed, and the original value from the preserved model is retained in the merged model.
-Appendix D Performance Drop Rate (PDR) for Adversarial Robustness
-The adversarial robustness is evaluated using the Performance Drop Rate (PDR) (Zhu et al., 2023), which is defined as follows:
-$\text{PDR}=\frac{\text{Metric}_{\text{no attack}}-\text{Metric}_{\text{attack}}}{\text{Metric}_{\text{no attack}}},$ (8)
-where $\text{Metric}_{\text{no attack}}$ denotes the performance metric without any prompt attack, and $\text{Metric}_{\text{attack}}$ represents the performance metric under the prompt attack.
-A smaller PDR indicates stronger adversarial defense against prompt attacks, implying better adversarial robustness.
-Appendix E Detailed Introduction to DARE
-DARE (Drop And REscale) (Yu et al., 2024) is a model sparsification method designed to reduce the redundancy of delta parameters in fine-tuned models while preserving their task-specific capabilities.
+Appendix H Detailed Introduction to DARE
+DARE (Drop And REscale) (Yu et al.
+2024) is a model sparsification method designed to reduce the redundancy of delta parameters in fine-tuned models while preserving their task-specific capabilities.
 In SFT, model parameters are optimized to unlock abilities for specific tasks, with the difference between fine-tuned and pre-trained parameters referred to as delta parameters.
 However, studies have shown that delta parameters are often highly redundant.
-DARE addresses this redundancy by randomly dropping a proportion $p$ of delta parameters (referred to as the drop rate) and rescaling the remaining ones by a factor of $1/(1-p)$.
+DARE addresses this redundancy by randomly dropping a proportion $p$ of delta parameters (referred to as the drop rate) and rescaling the remaining ones by a factor of 1/(1p)1/(1-p).
 This simple yet effective approach enables DARE to eliminate up to 99% of delta parameters with minimal impact on model performance, particularly in large-scale models, and it can be applied using only CPUs.
@@
 Experiments on AlpacaEval, GSM8K, and MBPP reveal that the merged LM has the potential to outperform any individual source LM, presenting a significant new discovery.
-Notably, the 7B model obtained through DARE merging, SuperMario v2, ranks first among models of the same scale on the Open LLM Leaderboard (Beeching et al., 2023).
+Notably, the 7B model obtained through DARE merging, SuperMario v2, ranks first among models of the same scale on the Open LLM Leaderboard (Beeching et al.
+2023).
 These improvements were achieved without the need for retraining, positioning DARE as an efficient and resource-friendly solution for model merging.
-Appendix F Detailed M3 and DARE Comparison Results
-Merging Methods Models Use M3 (Ours) Use DARE Instruction Following Mathematical Reasoning Code Generating
-AlpacaEval GSM8K MATH HumanEval MBPP
-Average Merging LM & Math Yes No 44.40 66.26 13.80 - -
-No Yes 44.22 66.57 12.96 - -
-Yes Yes 43.53 66.57 14.12 - -
-LM & Code Yes No 43.91 - - 37.20 34.40
-No Yes 38.81 - - 31.71 32.40
-Yes Yes 40.31 - - 36.59 37.00
-Math & Code Yes No - 63.61 14.02 8.54 19.20$\uparrow$
-No Yes - 56.18 10.28 6.10 7.80
-Yes Yes - 64.97 13.54 9.76 21.20
-Task Arithmetic LM & Math Yes No 41.65 66.34 13.74 - -
-No Yes 49.00 66.64 13.02 - -
-Yes Yes 44.90 67.32 13.74 - -
-LM & Code Yes No 46.64 - - 35.37 33.80
-No Yes 41.47 - - 35.98 33.00
-Yes Yes 45.20 - - 35.98 35.20
-Math & Code Yes No - 63.53 13.94 7.93 19.00
-No Yes - 65.05 13.96 10.37 9.80
-Yes Yes - 65.13 14.32 8.54 18.00
-TIES- Merging LM & Math Yes No 38.73 18.57 2.48 - -
-No Yes 37.92 18.04 2.34 - -
-Yes Yes 39.93 19.26 2.82 - -
-LM & Code Yes No 44.96 - - 25.61 30.80
-No Yes 43.13 - - 0.0 0.0
-Yes Yes 45.65 - - 26.83 33.20
-Math & Code Yes No - 64.75 14.16 9.76 21.4
-No Yes - 64.82 13.88 10.37 23.60
-Yes Yes - 64.75 14.78 9.15 19.60
-Table 8: Comparison of our method M3 and DARE. The best and second-best results are marked in bold and underlined fonts.
-We compare the two plug-and-play methods, M3 and DARE, by integrating each with three existing model merging techniques—Average Merging, Task Arithmetic, and TIES-Merging.
-The detailed experimental results are presented in Table 8.
-Appendix G Comparison of M3 and Existing Model Merging Approaches
-Models Merging Methods Instruction Following Mathematical Reasoning Code Generating Avg.
-AlpacaEval GSM8K MBPP
-LM & Math Average Merging 45.28 66.34 - 55.81
-Task Arithmetic 45.78 66.34 - 56.06
-TIES-Merging 38.63 14.56 - 26.60
-DARE 49.00 66.64 - 57.82
-DELLA 44.03 48.45 - 46.24
-M3 (Ours) 44.90 67.32 - 56.11
-LM & Code Average Merging 36.60 - 32.00 34.3
-Task Arithmetic 44.64 - 33.60 39.12
-TIES-Merging 41.85 - 0.0 20.93
-DARE 41.47 - 33.00 37.24
-DELLA 40.99 - 35.00 38.00
-M3 (Ours) 45.20 - 35.20 40.2
-Math & Code Average Merging - 56.17 8.20 32.19
-Task Arithmetic - 64.67 8.60 36.64
-TIES-Merging - 64.67 22.60 43.64
-DARE - 65.05 9.80 37.43
-DELLA - 63.08 20.2 41.64
-M3 (Ours) - 64.75 21.4 43.10
-Table 9: Comparison of our method M3 and existing model merging approaches. The best and second-best results are marked in bold and underlined fonts.
-For each baseline method—Average Merging, Task Arithmetic, TIES-Merging, and DARE—we combine it with our proposed M3 method and select the best-performing combination for each merged model (LM & Math, LM & Code and Math & Code).
-These optimal results are reported as M3 (Ours) in Table 9.
-As presented in Table 9, our proposed method M3 achieves the best overall performance across the three evaluated tasks.
-Importantly, M3 attains these improvements without involving complex architectural designs or intensive computations; it merely adjusts the contribution ratios of constituent models to realize substantial gains, thereby outperforming the DELLA approach.
-DELLA advances model sparsification techniques by building upon TIES-Merging and DARE, introducing a novel pruning strategy termed MAGPRUNE, which selects delta parameters based on their magnitudes.
-This strategy has inspired us to investigate new directions for enhancing the performance of merged models.
+Appendix I Integrating M3 into the TIES-Merging Model Merging Method
+Figure 5 shows the specific implementation approach to incorporating M3 into TIES-Merging.
+After the steps of trimming parameters with lower magnitudes and resolving sign disagreements, the two models to be merged are denoted as $M_{1}$ and $M_{2}$.
+During the M3 process, only the parameters that are preserved in both $M_{1}$ and $M_{2}$ are interpolated according to the model merging hyperparameter $\lambda_{m}$ to obtain the merged parameters.
+For parameters that are preserved in only one of the models, no interpolation is performed, and the original value from the preserved model is retained in the merged model.